### v2

# Basic representations of genus zero nonabelian Hodge spaces

Abstract.
In some previous work, we defined an invariant of genus zero nonabelian Hodge spaces taking the form of a diagram. Here, enriching the diagram by fission data to obtain a refined invariant, the enriched tree, including a partition of the core diagram into $k$ subsets, we show that this invariant contains sufficient information to reconstruct $k$+1 different classes of admissible deformations of wild Riemann surfaces, that are all representations of one single nonabelian Hodge space, so that the isomonodromy systems defined by these representations are expected to be isomorphic. This partially generalises to the case of arbitrary singularity data the picture of the simply-laced case featuring a diagram with a complete $k$-partite core. We illustrate this framework by discussing different Lax representations for Painlevé equations.
1. Introduction
1.1. General motivation: classification of (wild) nonabelian Hodge spaces
The main aim of this work is to formulate a convenient combinatorial way of identifying many conjectural isomorphisms between different genus zero nonabelian Hodge spaces, which should in turn extend to isomorphisms between the related isomonodromy systems, generalising known (multi)dualities between different Lax representations of Painlevé-type equations to the case of arbitrary irregular types.
The term ‘nonabelian Hodge space’, introduced in [10, Def. 7], refers to a class of manifolds that can be seen as moduli spaces for different types of objects. On one hand, they are moduli spaces of meromorphic connections with possibly irregular singularities on curves (the de Rham side). Via the Riemann-Hilbert-Birkhoff correpondence, they are also moduli spaces of generalised monodromy data, known as wild character varieties [8, 13] (the Betti side), and, via the nonabelian Hodge correspondence [4], as moduli spaces of irregular Higgs bundles (the Dolbeault side). Furthermore, they carry rich (symplectic, hyperkähler) geometric structures, and give rise to (isospectral and isomonodromic) integrable systems, including many classical integrable systems of interest in mathematical physics such as all Painlevé equations (see e.g. [10] for an overview).
A nonabelian Hodge space $\mathcal{M}$ depends on the choice of a tuple $(\Sigma,\mathbf{a},\bm{\Theta},\bm{\mathcal{C}})$, that we call a wild Riemann surface with boundary data. Here the triple $(\Sigma,\mathbf{a},\bm{\Theta})$ is a wild Riemann surface [8, 15], meaning that $\Sigma$ is a compact Riemann surface, $\mathbf{a}=\{a_{1},\dots,a_{m}\}\subset\Sigma$ is a finite set of points on $\Sigma$ corresponding to the singularities of the connections giving rise to points of $\mathcal{M}_{DR}$, the global irregular class $\bm{\Theta}$ encodes the type of singularities that these connections have at each singular point, and $\bm{\mathcal{C}}$ is a collection of conjugacy classes, encoding the formal monodromies of the connections.
Remarkably, it turns out that there exist some isomorphisms between moduli spaces defined by different wild Riemann surfaces with boundary data, corresponding to connections with different ranks, number of singularities and pole orders. While for most known examples such isomorphisms have not been worked out on all the three sides, it is expected that when this happens the moduli spaces are really isomorphic as nonabelian Hodge spaces, i.e. all the structures of the nonabelian Hodge package match in the natural way. In other words, we would like to view those different wild Riemann surfaces with boundary data as different realizations, or representations, of one single abstract nonabelian Hodge space $\mathcal{M}$, which raises the question of their classification. Notice that here we will only be considering representations coming from the case of vector bundles, however wild nonabelian Hodge spaces arising from vector bundles may also be isomorphic to nonabelian Hodge spaces arising from principal $G$-bundles, for $G$ a complex reductive group different from $\operatorname{GL}_{n}(\mathbb{C})$, see e.g. [12].
Varying the wild Riemann surface, one can consider moduli spaces of irregular connections in families, giving rise to isomonodromic deformations [8]. If the connections are written down explicitly in coordinates, this gives systems of nonlinear PDEs which include, among the simplest nontrivial examples, all Painlevé equations [33]. In this picture, the wild Riemann surface $(\Sigma,\mathbf{a},\bm{\Theta})$ giving rise to a given isomonodromy system corresponds to what is called a Lax representation, Lax pair, or linearisation for the system in the literature on integrable systems, albeit in a more abstract sense than is often the case (typically the term Lax pair is used to refer to an explicit parametrisation in coordinates of the coefficients of the connections). The existence of nontrivial isomorphisms between different nonabelian Hodge spaces is related to the existence of different Lax representations for many Painlevé-type equations, and the question of the classification of nonabelian Hodge spaces extends to the problem of finding all possible Lax representations for a given isomonodromy system.
1.2. Diagrams and isomorphisms between moduli spaces of connections
Interestingly, it turns out that in many cases, there is a clean combinatorial way to pass between different representations.
Possibly the simplest instance where such “dualities” between different moduli spaces of connections occur in genus zero is Harnad duality [1, 28, 47, 2], which relates dual connections on trivial vector bundles over the affine line of the form
$d-(A+P(Y-z)^{-1}Q)dz,\qquad\text{and}\qquad d+(Y+Q(A-z)^{-1}P)dz,$
where $P:W\to V$ and $Q:V\to W$ are linear maps, where $V,W$ are finite dimensional complex vector spaces, and $A\in\operatorname{End}(V)$, $Y\in\operatorname{End}(W)$. The connections on both sides have an irregular singularity at infinity, with a pole of order two, as well as regular singularities at finite distance, whose positions are given by the eigenvalues of $A$ and $Y$. Harnad showed that the isomonodromic deformations of these two connections, governed by the JMMS equations [32], are equivalent. The duality induces an isomorphism between the de Rham moduli spaces corresponding to both sides.
Harnad duality can be conveniently described in a graphical way as follows. The linear maps $A,Y,P,Q$ can be viewed as representations of the quiver drawn below, and the duality corresponds to exchanging the roles of the two vertices.
$V$$W$$A$$Y$$P$$Q$
When $A$ and $Y$ are semisimple, the vector spaces $V$ and $W$ can further be split into the corresponding eigenspaces, giving rise to a complete bipartite graph, and Harnad duality amounts to exchanging the two subsets of vertices.
This duality was enlarged by Boalch [5, 7, 9] in the simply-laced case, which is the case of connections on $\mathbb{P}^{1}$ having one irregular singularity, located at infinity, with an (unramified) pole of order at most 3, possibly together with regular singularities at finite distance: now there is a “multiduality” between more than two moduli spaces of connections.
This multiduality can also be conveniently encoded in a graphical way, in terms of a supernova quiver: it consists of a core, which is a complete $k$-partite graph for some integer $k$ (i.e. the core is a disjoint union of $k$ subsets of vertices, and each vertex in one subset is linked by one edge to every vertex belonging to a different subset) to which are then glued legs, each leg being a linear (i.e. type A) quiver.
The main statement is then as follows [7]: such a supernova quiver, together with some extra data, canonically determines $k$+1 moduli spaces of connections of simply-laced type, with in general different ranks and number of singularities, all isomorphic to each other. In each representation, the vertices of the core diagram are in one-to-one correspondence with the exponential factors of the connections. There is one generic representation, of maximal rank, with only one (irregular) singularity, located at infinity, and every $i=1,\dots,k$, defines a nongeneric representation, in which the vertices of the $i$-th subset of the core diagram correspond to regular singularities at finite distance, while all other core vertices correspond to exponential factors at infinity. In other words, one can read from the diagram $k$+1 representations of the same de Rham moduli space (see the example on the figure below). More concretely, the generic representation corresponds to connections of the form
$d-(Az+B)dz,$
with $A,B$ constant matrices, and $A$ semisimple, with $k$ distinct eigenvalues in one-to-one correspondence with the subsets of the partition of the core diagram. Furthermore, the equations governing the isomonodromic deformations of each representation are equivalent, so we can say that the different representations correspond to isomorphisms of isomonodromy systems.
Figure 1. Example of different readings of a (core) diagram in the simply-laced case. On the left, a complete 3-partite graph, with the different colours of the vertices corresponding to the partition. The three other figures on the right correspond to the nongeneric readings, each reading consisting in singling out one of the susbsets of the partition.
1.3. Basic operations on connections in genus zero and weak representations
From a more abstract point of view, Harnad duality is underlied by the Fourier-Laplace transform [28, 5, 7, 36]. It is an operation on (almost all) irreducible algebraic connections on Zariski open subsets of the affine line induced by the automorphism of the Weyl algebra of differential operators $A_{1}:=\mathbb{C}[z,\partial_{z}]$ defined by $z\mapsto-\partial_{z}$, $\partial_{z}\mapsto z$. Applied to such a connection, it gives another connection with in general different rank, number of singularities and pole orders.
Similarly, underlying the enlarged symmetry of the simply-laced case is the fact that the Fourier transform is actually part of a larger group of transformations acting on connections on the affine line: indeed, any matrix
$A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$
in $\mathrm{SL}_{2}(\mathbb{C})$ defines an automorphism of the Weyl algebra by $z\mapsto az+b$, $\partial_{z}\mapsto cz+d$. This induces an action of the group $\mathrm{SL}_{2}(\mathbb{C})$ on the set of isomorphism classes of irreducible algebraic connections on Zariski open subsets of the affine line $\mathbb{C}$ (excluding rank 1 connections with only a pole at infinity of order less than 2). The property of being of simply-laced type is preserved by this action, and the $k$+1 readings of the supernova quiver in the simply-laced case correspond to the different singularity patterns that appear in a given orbit.
More generally, there is a larger class of basic operations on irreducible irregular connections on $\mathbb{P}^{1}$, consisting of:
  • $\bullet$
    Möbius transformations of $\mathbb{P}^{1}$;
  • $\bullet$
    Twists by rank one connections;
  • $\bullet$
    $\operatorname{SL}_{2}(\mathbb{C})$ transformations.
We will often refer to the $\operatorname{SL}_{2}(\mathbb{C})$ transformations as symplectic transformations to avoid possible confusions with Möbius transformations. These operations play an essential role in the extension due to Deligne-Arinkin [17, 3] of the Katz algorithm for rigid local systems to the case of irregular connections, which provides a way to reduce any irreducible rigid connection on $\mathbb{P}^{1}$ to the trivial rank one connection by repeated application of such operations.
Möbius transformations and twists induce in a straightforward way isomorphisms of nonabelian Hodge spaces, in particular they do not change their Betti side description as wild character varieties. On the other hand, the Fourier-Laplace transform is expected to induce nontrivial isomorphisms of wild character varieties. This has been proven in some cases [45] but more work needs to be done for the general case, due to the fact that determining the Stokes data of the Fourier transform of a genus zero connection is quite involved [37, 29, 43, 23, 31, 38, 39, 19]. For this reason, since in the meantime we still want to view genus zero wild Riemann surfaces with boundary data related by basic operations as representations of the same nonabelian Hodge space, relying on the fact that basic operations on connections induce well-defined counterparts at the level of the corresponding wild Riemann surfaces with boundary data, we make the following definition:
Definition 1.1.
Let $\mathcal{M}=\mathcal{M}(\bm{\Sigma_{0}})$ be a genus zero nonabelian Hodge space, with $\bm{\Sigma_{0}}$ a wild Riemann surface with boundary data on $\mathbb{P}^{1}$. A weak representation of $\mathcal{M}$ is a wild Riemann surface with boundary data $\bm{\Sigma}$ that can be obtained from $\bm{\Sigma_{0}}$ by successive application of basic operations or admissible deformations.
In the rest of the article, we will allow ourselves to simply use the term representation to refer to weak representations. Given a (weak) representation $\mathbf{\Sigma}$ of a nonabelian Hodge space $\mathcal{M}$, we say that the representations obtained from $\Sigma$ by the action of a symplectic transformation defined by an element of $\operatorname{SL}_{2}(\mathbb{C})$, are its nearby representations. Finally, we will call basic representations of a nonabelian Hodge space the nearby representations of a representation of minimal rank.
1.4. Main results: from enriched trees to classes of nearby representations
The main purpose of this article is to develop a combinatorial way to find many weak representations of any nonabelian Hodge space in genus zero. Namely, we generalise the picture of obtaining different representations of a nonabelian Hodge space via different readings of a diagram from the simply-laced case to the case of genus zero nonabelian Hodge spaces with arbitrary singularity data. More specifically, we define some combinatorial data generalising the supernova quivers of the simply-laced case, from which several (expected to be) isomorphic nonabelian Hodge spaces can be explicitly obtained, corresponding to the different types of singularities appearing in an arbitrary $\operatorname{SL}_{2}(\mathbb{C})$ orbit, i.e. to different types of nearby representations of a given wild nonabelian Hodge space.
We build on our previous work [18], in which a construction of a diagram associated to an irregular connection on a Zariski open subset of $\mathbb{P}^{1}$ was formulated, for the general case of arbitrary irregular types. Furthermore it was shown that the dimension of the wild character variety is given by a formula involving the Cartan matrix of the diagram, in agreement with the philosophy of ‘Global Lie theory’ advocated in [9] suggesting to view nonabelian Hodge spaces as some global analogues of Lie groups. The main property of the diagram is that it is invariant under the symplectic $\operatorname{SL}_{2}(\mathbb{C})$ action, so that connections with different types of singularities have the same diagram. It turns out that, unlike for the simply-laced case, the diagram is not quite enough to reconstruct the corresponding nonabelian Hodge spaces, so some extra data will be needed. In particular, the core diagram is not a complete $k$-partite graph any more in the general case, and these extra data will include a canonical partition of the core diagram.
To obtain the appropriate extra data, as well as to give a precise meaning to the idea of ‘different types’ of singularity data, as was already sketched in [18, §5.3], we rely on our recent study of admissible deformations of wild Riemann surfaces [21, 20, 11, 22]. Being admissible deformations of each other defines an equivalence relation on the set of all possible wild Riemann surfaces of such connections, such that up to isomorphism the wild character variety $\mathcal{M}_{B}(\Sigma,\mathbf{a},\bm{\Theta},\bm{\mathcal{C}})$ is invariant under admissible deformations. This motivates the following definition:
Definition 1.2.
Let $\mathcal{M}$ be a nonabelian Hodge space. A class of representations of $\mathcal{M}$ is an equivalence class of admissible deformations of weak representations of $\mathcal{M}$.
In [11], we showed that a class of admissible deformations of wild Riemann surfaces is completely characterised by a pair $(g,\mathbf{F})$, where $g$ is the genus of the underlying Riemann surface $\Sigma$, $\mathbf{F}$ is a fission forest, i.e. a collection of fission trees (see [11, Def. 3.18]). We can thus write $\mathcal{M}=\mathcal{M}(g,\mathbf{F},\bm{\mathcal{C}})$ up to isomorphism, where $\bm{\mathcal{C}}$ is the datum of a conjugacy class in $\operatorname{GL}_{n}(\mathbb{C})$ for each leaf of some fission tree of $\mathbf{F}$ (with $n\geq 1$ an integer, the multiplicity of the leaf). Here we only consider the case $g=0$, so we will drop the genus from the notation and write $\mathcal{M}=\mathcal{M}(\mathbf{F},\bm{\mathcal{C}})$.
The data that will allow us to reconstruct the classes of nearby representations consist of an enriched tree, defined as follows:
Definition 1.3.
An enriched tree is a pair $\mathscr{T}=(\mathbf{T},\bm{\mathcal{C}})$, where
  • $\bullet$
    $\mathbf{T}$ is a short fission tree of , i.e. a fission tree such that all its vertices have height $\leq 2$.
  • $\bullet$
    If $N$ denotes the set of leaves of $\mathbf{T}$, $\bm{\mathcal{C}}$ is the datum of a conjugacy class $\mathcal{C}_{i}\in\operatorname{GL}_{n_{i}}(\mathbb{C})$, for some integer $n_{i}\geq 1$, for every $i\in N$.
We show that enriched trees provide an invariant of irregular connections on $\mathbb{P}^{1}$ under symplectic transformations:
Theorem 1.4 (Prop. 3.11, Def. 3.12, Cor. 3.13).
Any algebraic connection $(E,\nabla)$ on a Zariski open subset of $\mathbb{P}^{1}$ canonically determines an enriched tree $\mathscr{T}(E,\nabla)$, which depends only on its wild Riemann surface with boundary data. If $(E,\nabla)$ is irreducible, and is not a rank one connection with a pole of order less than 2 at infinity, and $A\in\operatorname{SL}_{2}(\mathbb{C})$, $A\cdot(E,\nabla)$ is well-defined and we have
$\mathscr{T}(A\cdot(E,\nabla))=\mathscr{T}(E,\nabla).$
Our main result is that this invariant is powerful enough to allow us to reconstruct all classes of nearby representations:
Theorem 1.5 (Cor. 3.10).
Let $\mathscr{T}=(\mathbf{T},\bm{\mathcal{C}})$ be an enriched tree, and $k$ be the number of principal subtrees of $\mathbf{T}$, i.e. the number of vertices of height 2. Assuming that the nonabelian Hodge space ($\mathcal{M}(\{\mathbf{T}\},\bm{\mathcal{C}})$ is nonempty, $\mathscr{T}$ canonically determines in an explicit way $k$ classes of representations $(\mathbf{F}_{i},\bm{\mathcal{C}_{i}})$, for $i=1,\dots,k$, of ($\mathcal{M}(\{\mathbf{T}\},\bm{\mathcal{C}})$, such that, if $(\bm{\Theta},\bm{\mathcal{C}})$ is any wild Riemann surface with boundary data on $\mathbb{P}^{1}$ with enriched tree $(\mathbf{T},\bm{\mathcal{C}})$, then ($(\{\mathbf{T}\},\bm{\mathcal{C}})$ and $(\mathbf{F}_{i},\bm{\mathcal{C}_{i}})$, for $i=1,\dots,k$ are all the classes of its nearby representations of ($\mathcal{M}(\{\mathbf{T}\},\bm{\mathcal{C}})$.
While we refer the reader to the body of the article for the detailed construction of the classes of representations, let us now briefly summarise their properties and discuss the link with the diagrams of [18]. Recall that if $(E,\nabla)$ is an algebraic connection on a Zariski open subset of $\mathbb{P}^{1}$, at each of its singularities it possesses a finite set of Stokes circles, which basically correspond to the different exponential factors $e^{q}$ appearing in the asymptotics of the horizontal sections of the connection close to the singularity. In our language they are encoded by the (modified) global irregular class $\bm{\breve{\Theta}}$ which is the collection of all Stokes circles of $(E,\nabla)$, with some integer multiplicities.
Theorem 1.6 (Prop. 3.14, 3.15).
[…]
Let $(E,\nabla)$ be an irreducible irregular connection on $\mathbb{P}^{1}$ with enriched tree $(\mathbf{T},\bm{\mathcal{C}})$, and $(E^{\prime},\nabla^{\prime})\in\operatorname{SL}_{2}(\mathbb{C})\cdot(E,\nabla)$ an element in its orbit. We have:
  1. (1)
    There is a bijection between $N$ and the set of Stokes circles of $(E^{\prime},\nabla^{\prime})$. Furthermore:
    • $\bullet$
      If $(E^{\prime},\nabla^{\prime})$ is in the generic class of representations, all its Stokes circles are of generic form, i.e. are at infinity and are of slope $\leq 2$.
    • $\bullet$
      Otherwise, if $(E^{\prime},\nabla^{\prime})$ is in the $i$-th nongeneric class of representations for $i=1,\dots k$,
      • the Stokes circles corresponding to elements of $N_{i}^{-}$ via the bijection are at finite distance (i.e. different from $\infty\in\mathbb{P}^{1}$),
      • the Stokes circles corresponding to elements of $N_{i}^{+}$ are at infinity and of slope $>2$,
      • the Stokes circles corresponding to $N_{j}$ for every $j\neq i$ are of generic form.
  2. (2)
    If $(E^{\prime},\nabla^{\prime})$ is in the $i$-th nongeneric class of representations, the number of its singularities at finite distance is equal to the number of principal subtrees of $\mathbf{T}_{i}$ with root at height 1.
  3. (3)
[…]
    • $\bullet$
      The generic class of representations has rank
      $r_{gen}=\sum_{i=1}^{k}\left(\sum_{I\in N_{i}}n_{I}r_{I}\right).$
    • $\bullet$
      The $i$-th nongeneric class of representations has rank
      $r_{i}=\sum_{I\in N_{i}^{+}}n_{I}R_{I}+\sum_{\begin{subarray}{c}l=1\\ l\neq i\end{subarray}}^{k}\left(\sum_{I\in N_{l}}n_{I}r_{I}\right).$
The link with the diagrams of [18] is as follows. Any irregular connection on $\mathbb{P}^{1}$ canonically determines both a diagram $\Gamma$ and an enriched tree $\mathscr{T}$, both invariant under symplectic transformations. Our main result says that the datum of the enriched tree allows us to construct $k+1$ different readings of the diagram, corresponding to the classes of representations in an $\operatorname{SL}_{2}(\mathbb{C})$-orbit. The enriched tree is thus a finer invariant that contains in general strictly more information than the diagram, i.e. we can view it as an enriched diagram. Concretely, the enriched tree determines explicitly the diagram, as well as a partition of the set of its core vertices into $k$-subsets.
Notice that, from the perspective of our general framework, the simply-laced case of [7] corresponds to the situation where all vertices of the generic fission tree have integer heights (belonging to the set $\{1,2\}$). In this case, the datum of the short fission tree is equivalent to the datum of the complete $k$-partite core diagram, which is consistent with the main result of [7] saying that everything can be reconstructed from the supernova quiver.
Let us also mention that the fission trees of [11] are closely related to combinatorial objects appearing in the study of the singularities of plane algebraic curves: they can be seen as a variant of the Eggers-Wall tree [46]. In some sense, we can thus view the short fission tree as a global analogue of Eggers-Wall trees, for the case of several singularities located at different points on $\mathbb{P}^{1}$.
We apply our framework to the case of nonabelian Hodge spaces related to Painlevé equations, which have complex dimension 2 and constitute the simplest nontrivial examples of nontrivial nonabelian Hodge spaces. Concretely for all Painlevé moduli spaces we describe their classes of basic representations, which are classes of nearby representations of the standard rank 2 Lax representations known since the early 20th century [26, 27, 33]. In this way we recover many known alternative Lax representations as basic representations different from the standard one, and we can view the remaining basic representations as providing new Lax representations (in our abstract sense).
While we refer the reader to the last section of the article for many more pictures of fission trees and basic representations for Painlevé moduli spaces, a concrete example is already summarized on Figure LABEL:fig:PIII_tree_and_diagram, which corresponds to the classes of basic representations of the Painlevé III moduli space which are the classes of nearby representations of the standard rank 2 Painlevé III Lax representation (see §4 for a more detailed discussion of this example). The figure illustrates how, starting from the short fission tree, the other classes of representations can be obtained, yielding different readings of the diagram.
To conclude this introduction, let us sketch briefly sketch the main ideas leading to these results. An essential observation, already made in [7] in the simply-laced case, and in [18] in the general case, is that one can associate to any Stokes circle $I$ an element $\lambda(I)\in\mathbb{P}^{1}$ that we will call its Fourier sphere coefficient, such that the $\operatorname{SL}_{2}(\mathbb{C})$ action on the set of all possible Stokes circles (induced by the action on connections via the stationary phase formula) behaves in a very nice way with respect to those: it acts by homographies on the Fourier sphere $\mathbf{P}^{1}$, the copy of $\mathbb{P}^{1}$ where those coefficients live. Concretely, for
$A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{C}),$
letting $h_{A}$ denote the corresponding homography
$h_{A}:z\mapsto\frac{az+b}{cz+d},$
we have that $A$ acts on the Fourier sphere coefficients by $h_{A}$.
Then, for any connection $(E,\nabla)$ one can consider the set $\{\lambda_{1},\dots,\lambda_{k}\}$ of the distinct Fourier sphere coefficients of its Stokes circles, which defines the integer $k$.
Using the results of [11] on admissible deformations, it is then quite straightforward to see that up to admissible deformations, the irregular class of $A\cdot(E,\nabla)$ only depends on whether we have $h_{A}(\lambda_{i})=\infty$ for some $i=1,\dots,k$, and if so for which $i$, which leads to the nearby representations: the generic representation corresponds to the case where $h_{A}(\lambda_{i})\neq\infty$ for all $i=1,\dots k$, and the $i$-th nongeneric representation to the case where $h_{A}(\lambda_{i})=\infty$. Each distinct Fourier sphere coefficient $\lambda_{i}$ corresponds to one principal subtree of the fission tree of the generic representation.
1.5. Outlook
An interesting question for future work could be to investigate to which extent one can devise a general procedure to construct, for any such abstract Lax representation, a corresponding fully explicit Lax representation, that is suitable explicit parametrisations of connections with the singularity data determined by the abstract Lax representation, such that their isomonodromic deformations imply that some function $y$ appearing in the parametrisation satisfies the corresponding Painlevé equation. Indeed, the search for alternative Lax representations for Painlevé-type equations has been an active topic of research [34, 35, 44, 41, 16].
Furthermore, a natural extension of this work is the question of actually finding explicitly the isomorphisms expected to exist between the nonabelian Hodge spaces corresponding to the different representations, and of understanding how the spaces of isomonodromic times for different representations are related, so as to obtain, as in the simply-laced case, a full correspondence between the isomonodromy systems.
1.6. Contents
The text is organised as follows. In section 2, we review the necessary facts about singularity data of irregular connections, admissible deformations, fission forests, and diagrams, and describe how these data are transformed under symplectic transformations, relying on the results of [18] and slightly extending them to deal with admissible deformations. In section 3, we define the generic form of a Stokes circle and a modified irregular class, Fourier sphere coefficients, and establish our main results, in particular we provide explicit formulas for the rank of all nearby representations. Finally, in section 4, we apply this framework to describe in detail the basic representations of all Painlevé moduli spaces.
Acknowledgements
I am funded by the PNRR Grant CF 44/14.11.2022, “Cohomological Hall Algebras of Smooth Surfaces and Applications”, led by Olivier Schiffmann. I also thank the funding of FCiências.ID, the great working environment of the Grupo de Física Matemática, where most of this work was carried out, and Philip Boalch, Gabriele Rembado, Giordano Cotti, Giulio Ruzza, Davide Masoero, Gabriele Degano, Julian Barragán Amado, for useful discussions.
2. Symplectic action on fission data of irregular connections on $\mathbb{P}^{1}$
In this section, we briefly review the necessary facts about formal data of connections with irregular singularities on the Riemann sphere, their transformation under Fourier transform and more general symplectic transformations, admissible deformations, and combinatorial objects (fission forests, diagrams) encoding in a useful way parts of the information contained in the formal data.
2.1. Formal data of irregular connections
Let us first recall the description of formal data of irregular connections on the Riemann sphere. The Turritin-Levelt theorem states that any connection on the formal punctured disc can be decomposed as a direct sum of elementary connections having only one exponential factor. This can be formulated in different ways; the one we will be using involves viewing exponential factors at a singularity as sections of a local system over the circle of directions around this singularity (see e.g. [18, 15] for more details).
2.1.1. The exponential local system
Let $\Sigma=\mathbb{P}^{1}$ and $a\in\Sigma$. Let $\pi_{a}:\widehat{\Sigma}_{a}\to\Sigma$ the real oriented blow-up at $a$ of $\Sigma$. The preimage $\partial_{a}:=\pi^{-1}(a)$ is a circle whose points correspond to the directions around $a$. An open subset of $\partial_{a}$ corresponds to a germ of sector at $a$.
Let $z_{a}$ be a local complex coordinate on $\Sigma$ in a neighbourhood of $a$, vanishing at $a$. The exponential local system $\mathcal{I}_{a}$ is a local system of sets (i.e. a covering space) over $\partial_{a}$ whose sections are germs of holomorphic functions on sectors of the form
$\sum_{i}b_{i}z_{a}^{-k_{i}},$
where $k_{i}\in\mathbb{Q}_{>0}$, and $b_{i}\in\mathbb{C}$. The connected component of such a local section is a finite order cover of the circle $\partial$. More precisely, let $r$ the smallest integer such that the expression $q=\sum_{i}b_{i}z_{a}^{-k_{i}}$ is a polynomial in $z_{a}^{-1/r}$. The corresponding holomorphic function is multivalued, and becomes single-valued when passing to a finite cover $t_{a}^{r}=z_{a}$. Therefore, the corresponding connected component, which we denote by $\langle q\rangle$, is a $r$-sheeted cover of $\partial_{a}$. As a topological space, it is homeomorphic to a circle, and $\mathcal{I}_{a}$ is thus a disjoint union of (an infinite number of) such Stokes circles.
The integer $r=:\operatorname{ram}(q)$ is the ramification order of the circle $I=\langle q\rangle$. If $r=1$ we say that the circle $\langle q\rangle$ is unramified.
If $d\in\partial_{a}$, we denote by $(\mathcal{I}_{a})_{d}$ the fibre of $\mathcal{I}$ over the direction $d$. Taking $d$ as a basis point of $\partial_{a}$, the monodromy of $\mathcal{I}_{a}$ is the automorphism
$\rho_{a}:(\mathcal{I}_{a})_{d}\to(\mathcal{I}_{a})_{d}$
of the fibre $(\mathcal{I}_{a})_{d}$ obtained when going once around $\partial_{a}$ in the positive direction.
Finally we define the global exponential local system as the disjoint union
$\mathcal{I}:=\sqcup_{a\in\Sigma}\mathcal{I}_{a}.$
It naturally comes with a projection $\pi:\mathcal{I}\to\partial$, where
$\partial:=\sqcup_{a\in\Sigma}\partial_{a}.$
If $I\subset\mathcal{I}_{a}$ is a Stokes circle at $a$ for some $a\in\mathbb{P}^{1}$, by a slight abuse of notation we will write $\pi(I):=a$.
2.1.2. Local and global formal data
Definition 2.1.
An irregular class at $a$ is a function $\Theta_{a}:\pi_{0}(\mathcal{I}_{a})\to\mathbb{N}$ with compact support. The rank of an irregular class $\Theta_{a}$ is
[…]
where $n_{I}=\Theta_{a}(I)$ is the multiplicity of the circle $I$. If $n_{I}\neq 0$ we say that $I$ is an active circle of $\Theta_{a}$, or a Stokes circle of $\Theta_{a}$.
The Turritin-Levelt theorem says that any rank $r$ connection on the formal punctured disc at $a$ uniquely determines a pair $(\Theta_{a},\mathcal{C}_{a})$, where $\Theta_{a}$ is a rank $r$ irregular class, and $\mathcal{C}_{a}$ is the datum of a conjugacy class $\mathcal{C}_{I}\in GL_{n_{I}}(\mathbb{C})$ for any active circle $I$ of $\Theta_{a}$ with multiplicity $n_{I}$. The pair $(\Theta_{a},\mathcal{C}_{a})$ constitutes the formal data of the connection and determines it completely up to isomorphism.
Passing to the global situation, to any algebraic connection $(E,\nabla)$ on a Zariski open subset $\Sigma\smallsetminus\mathbf{a}$ of $\Sigma$, by restricting the connection on a small disc around each singularity and considering its formalization there, one associates global formal data $(\bm{\Theta},\bm{\mathcal{C}}):=(\Theta_{a},\mathcal{C}_{a})_{a\in\mathbf{a}}$.
The Stokes circles of $\Theta_{a}$ at each singularity $a\in\Sigma$ correspond to the exponential factors of the connection in the Turritin-Levelt normal form at $a$, and the conjugacy classes $\mathcal{C}_{a}$ encode its the formal monodromy. In particular, the connection has a regular singularity at $a$ if and only if its only Stokes circle at $a$ is the tame circle $\langle 0\rangle_{a}$.
Remark 2.1.
Notice that if we wanted to work with the full nonabelian Hodge package, we would need to consider weighted conjugacy classes (cf. [6]). Here for simplicity, since we won’t be trying in this work to prove that the different classes of representations correspond as one expects to isomorphisms between the full nonabelian Hodge spaces, we will just work will standard (unweighted) conjugacy classes.
2.1.3. Modified formal data
It turns out that the action of the Fourier transform on formal data of connections is more easily described in terms of modified formal data, obtained as follows: if $i$ is a tame circle at finite distance ($\pi(i)\neq\infty$ and $q_{i}=0$) then replace $n_{i}$ by the integer
$m_{i}=\text{rank}(A-1),$
where $A\in\mathcal{C}_{i}$. Thus $A=1+vu$ for a surjective map $u:\mathbb{C}^{n_{i}}\to\mathbb{C}^{m_{i}}$ and an injective map $v:\mathbb{C}^{m_{i}}\to\mathbb{C}^{n_{i}}$. Then replace $\mathcal{C}_{i}$ by the conjugacy class $\breve{\mathcal{C}}_{i}$ of $(1+uv)$ in $\operatorname{GL}_{m_{i}}(\mathbb{C})$; this class is called the child of $\mathcal{C}_{i}$ in [9, appendix]. If $\pi(i)=\infty$ or if $i$ is not tame ($q_{i}\neq 0$) then we do no modification: $m_{i}=n_{i}$ and $\breve{\mathcal{C}}_{i}=\mathcal{C}_{i}$. This defines modified formal data $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$, where $\bm{\breve{\Theta}}$ is the collection of circles with multiplicities $m_{i}\geq 1$.
[…]
2.2. Admissible deformations, fission data, and forests
We now recall a few definitions and results about admissible deformations of irregular connections, involving the notions of levels and fission trees, and diagrams. We refer the reader to [11, 18] for more details.
2.2.1. Local version
First, to define the notion of admissible deformation of an irregular class, we need to recall the notion of (full) irregular type.
Definition 2.2.
A full irregular type at $a\in\Sigma$ of rank $n$ is a Galois-closed list $T=(q_{1},\dots,q_{n})$ of exponential factors in some direction $d\in\partial$ (i.e. $q_{i}$ is a section of $\mathcal{I}_{a}$ in a germ of sector around $d$).
Any full irregular type $T=(q_{1},\dots,q_{n})$ defines an associated irregular class $\Theta$ in the following way: its active circles are the connected components $\langle q_{i}\rangle$ in $\mathcal{I}$ of the exponential factors, and the multiplicity $n_{i}$ of $\langle q_{i}\rangle$ is the multiplicity of $q_{i}$ in the list $(q_{1},\dots,q_{n})$. Passing from a full irregular type to an irregular class amounts to forgetting the ordering of the exponential factors.
Definition 2.3.
Let $T=(q_{1},\dots,q_{n})$ and $T^{\prime}=(q^{\prime}_{1},\dots,q^{\prime}_{n})$ be rank $n$ full irregular types at $a$ in some direction $d$. We say that $T$ and $T^{\prime}$ are admissible deformations of each other, and denote this by $T\sim T^{\prime}$, if we have
$\operatorname{slope}(q_{i}-q_{j})=\operatorname{slope}(q^{\prime}_{i}-q^{\prime}_{j}),$
for all $i,j=1,\dots,n$.
If $\Theta$ and $\Theta^{\prime}$ are two rank $n$ irregular classes at $a\in\Sigma$, we say that they are admissible deformations of each other, and write $\Theta\sim\Theta^{\prime}$, if there exist two full irregular types $T$ and $T^{\prime}$ with respective associated irregular classes $\Theta$ and $\Theta^{\prime}$ such that $T\sim T^{\prime}$.
In particular, if $\Theta\sim\Theta^{\prime}$, denoting by $N$ and $N^{\prime}$ the sets of active circles of $\Theta$, $\Theta^{\prime}$ respectively, there exists a bijection $\phi:N\to N^{\prime}$ such that for each active circle $I\in N$, $\phi(I)$ is an admissible deformation of $I$, and $I$, $I^{\prime}$ have the same multiplicity. We say that such a bijection is compatible.
Definition 2.4.
Let $(\Theta,\mathcal{C})$ and $(\Theta^{\prime},\mathcal{C}^{\prime})$ two sets of local formal data. We say that they are admissible deformations of each other if $\Theta\sim\Theta^{\prime}$ and there exists a compatible bijection $\phi:N\to N^{\prime}$ between their respective sets of active circles $N,N^{\prime}$ such that $\mathcal{C}^{\prime}_{\phi(I)}=\mathcal{C}_{I}$ for all $I\in N$.
It turns out that there exist simple numerical criteria characterising whether two irregular classes are admissible deformations of each other.
Definition 2.5.
Let $\Theta$ be an irregular class at $a\in\Sigma$. The levels of $\Theta$ are the elements of the set
(2.1) $\mathrm{Levels}(\Theta)=\{\operatorname{slope}(q_{\alpha}-q_{\beta})\;|\;\alpha,\beta\in\mathtt{I}_{d}\}\setminus\{0\}\subset\mathbb{Q}_{>0}$
where $\mathtt{I}\subset\mathcal{I}$ is the finite subcover underlying $\Theta$ and $\mathtt{I}_{d}$ is any fibre of $\mathtt{I}$.
It is clear that two irregular classes which are admissible deformations of each other have the same levels.
Let us first consider the case of an irregular class having only one Stokes circle. Let $I$ be a Stokes circle at $a$. We can always write $I=\langle q\rangle_{a}$, with $q=\sum_{i=1}^{p}a_{i}z_{a}^{-k_{i}}$, for some integer $p$, with $k_{1}>\dots>k_{p}$ and $a_{i}\neq 0$ for every $i=1,\dots,p$. The rational numbers $k_{i}$ are uniquely defined by $I$, and we call them the exponents of $I$. We have the following:
Proposition 2.6.
a) Two Stokes circles $I,J\subset\mathcal{I}_{a}$ are admissible deformations of each other if and only if $\operatorname{Levels}(I)=\operatorname{Levels}(J)\subset\mathbb{Q}$.
b) A subset $(k_{1}>k_{2}>\cdots>k_{m})\subset\mathbb{Q}_{>0}$ is the set of levels of some circle $I\subset\mathcal{I}_{a}$ if and only if
(2.2) $k_{1},k_{2},\ldots,k_{m}$ have strictly increasing common denominators $>1$.
In other words if $d_{i}$ is the denominator of $k_{i}$ (in lowest terms) and
(2.3) $r_{i}\text{ is the lowest common multiple of }d_{1},d_{2},\ldots,d_{i}$
for each $i$ (so that $r_{i}\;|\;r_{i+1}$), then $1<r_{1}<r_{2}<\cdots<r_{m}$.
c) Let $I$ be a Stokes circle, and let $(k_{1}>\dots>k_{p})$ denote its set of exponents. Then $\operatorname{Levels}(I)$ is the largest possible subset of $\{k_{1},\dots,k_{p}\}$ of the form $k_{i_{1}}>\dots>k_{i_{m}}$ for with $1=i_{1}<\dots<i_{m}\leq p$ such that $k_{i_{1}},\dots,k_{i_{m}}$ have strictly increasing common denominators $>1$.
Passing to the general case with irregular classes having several Stokes circles, to characterise admissible deformations, we need more data than the levels of each individual Stokes circle. To this end we introduce the notion of common part and fission exponent of a pair of Stokes circles.
If $q=\sum_{i}b_{i}z_{a}^{-i/r}$ is an exponential factor at $a$, and $k\in\mathbb{Q}_{\geq 0}$, we write
$q_{\geq k}:=\sum_{i,i/r\geq k}b_{i}z^{-i/r}.$
In turn, if $I=\langle q\rangle$ is a Stokes circle, we set $I_{\geq k}:=\langle q_{\geq k}\rangle$. We define in a similar way truncations $I_{>k}$, $I_{\leq k}$, $I_{<k}$.
Definition 2.7.
Let $I,\widehat{I}$ be two Stokes circles at the same point with respective ramification orders $r,\widehat{r}$. The common part of $I$ and $\widehat{I}$ is the circle $I_{c}=I_{\geq k}=\widehat{I}_{\geq k}$, where
$k=\min\left(l\in\frac{1}{r\widehat{r}}\mathbb{Z}_{\geq 0}\;\middle|\;I_{\geq l}=\widehat{I}_{\geq l}\right).$
The fission exponent of $I$ and $I^{\prime}$ is the rational number
$f_{I,\widehat{I}}=\max(\operatorname{slope}(I_{<k}),\operatorname{slope}(\widehat{I}_{<k}))\in\frac{1}{r\widehat{r}}\mathbb{Q}.$
If $I=\widehat{I}$ then $I_{c}=I=\widehat{I}$ and $f_{I,\widehat{I}}=0$, otherwise $f_{I,\widehat{I}}>0$. $I_{c}=0$ if and only if the leading terms of $I$ and $\widehat{I}$ have different Galois orbits, in this case we say that $I$ and $\widehat{I}$ have no common part.
Lemma 2.8.
Let $I$ and $\widehat{I}$ be two Stokes circles with ramification orders $r,\widehat{r}$. The set of slopes among the rational numbers $\operatorname{slope}({q_{i}}-{\widehat{q}_{j}})$, for $i=0,\dots,r-1$, $j=0,\dots,\widehat{r}-1$ is of the form
$\operatorname{Levels}(I_{c})\sqcup\{f_{I,\widehat{I}}\}\subset\mathbb{Q}_{\geq 0},$
i.e. it consists of the levels of the common part of together with their fission exponent.
This motivates the following definition:
Definition 2.9.
A (local) “fission datum” is a pair $\mathcal{F}=(\mathcal{L},f)$ where $\mathcal{L}$ is a multiset
$\mathcal{L}=L_{1}+\cdots+L_{n}$
of level data and $f$, the fission exponents, are the choice of a rational number $f_{ij}=f_{ji}\in\mathbb{Q}_{\geq 0}$, for all $i,j\in\{1,\ldots,n\}$.
A (local) irregular class determines a fission datum as follows. If $\Theta=\sum_{1}^{n}I_{i}$ is a rank $n$ irregular class at some point $a$ (where the Stokes circles $I_{i}$ are not necessarily distinct), then define $L_{i}=L(I_{i})$ to be the level datum of $I_{i}$ for each $i$, and define $\mathcal{L}(\Theta):=\sum_{1}^{n}L_{i}$ to be the corresponding multiset of level data. Then by taking $f_{ij}=f_{I_{i},I_{j}}$ to be corresponding fission exponents, this determines the fission datum $\mathcal{F}(\Theta)=(\mathcal{L}(\Theta),f)$ of the irregular class $\Theta$. Note that the multiplicity of any given Stokes circle $I_{j}$ in the class $\Theta$ is determined from the fission data by the recipe:
$\Theta(I_{j})=\absolutevalue{\{i=1,2,\ldots,n\;|\;f_{ij}=0\}}.$ | .
The fission datum characterises classes of admissible deformations of (local) irregular classes.
Theorem 2.10.
Two (local) irregular classes $\Theta$ and $\Theta^{\prime}$ are admissible deformations of each other if and only if they have the same fission datum.
Notice that not any fission datum $\mathcal{F}$ actually comes from an irregular class $\Theta$. To characterise such fission data, as well as provide a more graphical way of encoding them, the notion of a (twisted) fission tree is introduced in [11]. We will not be needing to recall the exact axioms for fission trees here, so we refer the reader once more to loc. cit. for the full details. Let us just briefly summarize the construction of the fission tree $\mathcal{T}(\Theta)$ associated to an irregular class $\Theta$.
  • $\bullet$
    To each active circle $I$ of $\Theta$ with level datum $L$ one associates a full branch $T_{I}$, which is the datum of a copy of $\mathbb{R}_{\geq 0}$, and of a set of distinguished points on it, the vertices, defined as the set of all possible exponents of a Stokes circle with level datum $L$. The real coordinate on a full branch will be called the height.
  • $\bullet$
    The tree $\mathcal{T}(\Theta)$ is the topological space obtained by gluing for each pair of disctinct active circles $I,J$ the corresponding full branches together above a height defined as the smallest height of any vertex in some full branch strictly greater than the fission exponent $f_{I,J}$.
In brief, the leaves of the tree are in bijection with the active circles of $\Theta$, and the tree contains the information about the level data as well as the fission exponents. In particular the datum of the fission tree $\mathcal{T}(\Theta)$ is equivalent to the datum of $\mathcal{F}(\Theta)$.
Finally, if we include conjugacy classes of formal monodromies, a set of local formal data $(\Theta,\mathcal{C})$ determines a fission tree $\mathbf{T}:=\mathcal{T}(\Theta)$, and $\mathcal{C}$ amounts to the datum of a conjugacy class for each leaf of $\mathbf{T}$. Class of admissible deformations of local formal data is characterised by such pairs $(\mathbf{T},\mathcal{C})$.
2.2.2. Global version
Until now in this paragraph we have dealt only with the local situation. The global case is described in a similar way, by collecting together the local data at each singularity.
[…]
Corollary 2.11.
Two global irregular classes $\mathbb{P}^{1}$ are admissible deformations of each other if and only if they have the same global fission datum, or equivalently the same fission forest.
As a consequence, a class of admissible deformations of wild Riemann surfaces with boundary data is characterised by a pair $(\mathbf{F},\bm{\mathcal{C}})$, where $\mathbf{F}$ is a fission forest, and $\mathcal{C}$ the datum of a conjugacy class for each leaf of $\mathbf{F}$.
2.3. Diagrams
Another invariant that we can attach to any connection $(E,\nabla)$ on a Zariski open subset of $\mathbb{P}^{1}$ is a diagram $\Gamma(E,\nabla)$ [18]. As for the forest, the diagram only depends on the (modified) formal data $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ of the connection.
The diagram $\Gamma(E,\nabla)=\Gamma(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})=(N,B)$ of the connection $(E,\nabla)$ has the following structure: it consists of a core diagram $\Gamma_{c}(E,\nabla)=\Gamma_{c}(\bm{\breve{\Theta}})$ with nodes $N$, to which is then glued, onto each core node, a single leg determined by a conjugacy class.
The set of core nodes $N$ is the set of Stokes circles of the modified irregular class $\bm{\breve{\Theta}}$. In terms of the non-modified irregular class $\bm{\Theta}$, $N$ is the disjoint union of three finite sets:
$N=$$\{\text{Stokes circles of $(E,\nabla)$ at $\infty$}\}\ \cup$
$\{\text{wild Stokes circles of }(E,\nabla)\text{ not at $\infty$}\}\ \cup$
{tame Stokes circles of ($\{\text{tame Stokes circles of }(E,\nabla)\text{ not at $\infty$ with nontrivial formal monodromy}\}.$
The matrix of edge multiplicities between the core nodes is then defined as follows. For any Stokes circle $I=\langle q\rangle$, let $\alpha_{I}=\text{Irr}(I),\beta_{I}=\operatorname{ram}(I)\in\mathbb{N}$ denote its irregularity and ramification numbers, so that $I$ has slope $\alpha_{I}/\beta_{I}$. We define an integer $B^{\infty}_{IJ}$ as follows:
(2.4) $B^{\infty}_{IJ}$ $=A_{IJ}-\beta_{I}\beta_{J},\qquad\text{if $I\neq J$,}$
(2.5) $B^{\infty}_{II}$ $=A_{II}-\beta_{I}^{2}+1,\quad\,\text{otherwise,}$,
where $A_{IJ}=\operatorname{Irr}(\operatorname{Hom}(I,J))$ is the irregularity of the irregular class $\operatorname{Hom}(I,J)$.
Definition 2.12.
Suppose $I,J\in N$ are Stokes circles, and write $a_{I}=\pi(I),a_{J}=\pi(J)\in\mathbb{P}^{1}$ for the underlying points of the Riemann sphere. Define $B_{IJ}\in\mathbb{Z}$ as follows:
  1. (1)
    If $a_{I}=a_{J}=\infty$ then $B_{IJ}=B^{\infty}_{IJ}$,
  2. (2)
    If $a_{I}=a_{J}\neq\infty$ then $B_{IJ}=B_{JI}=B^{\infty}_{IJ}-\alpha_{I}\beta_{J}-\alpha_{J}\beta_{I}$,
  3. (3)
    If $a_{I}=\infty$ and $a_{J}\neq\infty$ then $B_{IJ}=\beta_{I}(\alpha_{J}+\beta_{J})$,
  4. (4)
    If $a_{I}\neq\infty$ and $a_{J}=\infty$ then $B_{IJ}=B_{JI}=\beta_{J}(\alpha_{I}+\beta_{I})$,
  5. (5)
    If $a_{I}\neq\infty,\;a_{J}\neq\infty$ and $a_{I}\neq a_{J}$ then $B_{IJ}=B_{JI}=0$.
This completes the definition of the core diagram. To obtain the full diagram, one glues to each core node $I\in N$ a leg defined by the conjugacy class $\breve{\mathcal{C}}_{i}$.
Remark 2.2.
The core diagram $\Gamma_{c}(\bm{\breve{\Theta}})$ is entirely determined by the fission data of $\bm{\breve{\Theta}}$ together with the slopes of all Stokes circles at finite distance, because these data determine the number $A_{IJ}=\operatorname{Irr}(\operatorname{Hom}(I,J))$. In general, the diagram contains strictly less information, because $A_{IJ}$ is a sum of different contributions, coming from the levels of the common parts of $I,J$ and the fission exponent $f_{I,J}$. As we shall review below, the interest of the (core) diagram is that, unlike the fission data, it is invariant under symplectic transformations
2.4. Fourier transform and symplectic transformations
We now describe how the $\operatorname{SL}_{2}(\mathbb{C})$ action on connections acts on their formal data and fission data.
2.4.1. Fourier transform
The Fourier transform is the automorphism of the Weyl algebra $A_{1}=\mathbb{C}[z]\langle\partial_{z}\rangle$ defined by $z\mapsto-\partial_{z},\;\partial_{z}\mapsto z$. It induces a transformation on modules on the Weyl algebra: the Fourier-Laplace transform $F\cdot M$ of a $\mathbb{C}[z]\langle\partial_{z}\rangle$-module $M$ is the $\mathbb{C}[z]\langle\partial_{z}\rangle$-module obtained from $M$ by setting: $z\mapsto-\partial_{z},\;\partial_{z}\mapsto z$. The Fourier transform induces a transformation on the set of isomorphism classes of irreducible connections (if we exclude the case of the trivial rank one connection on $\mathbb{P}^{1}$).
The stationary phase formula [24, 42] states that the modified formal data of the Fourier transform $(E^{\prime},\nabla^{\prime})=F\cdot(E,\nabla)$ are determined by the modified formal data of $(E,\nabla)$.
Theorem 2.13.
There exists a bijection, that we also denote by $F$, from the set of all modified formal data to itself, such that if $(E,\nabla)$ is a irreducible algebraic connection on a Zariski open subset of $\mathbb{P}^{1}$, $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ its modified formal data, $(E^{\prime},\nabla^{\prime})$ is the Fourier transform of $(E,\nabla)$ and $(\bm{\breve{\Theta}^{\prime}},\bm{\breve{\mathcal{C}}}^{\prime})$ its modified formal data, the following diagram commutes:
${(E,\nabla)}$${(E^{\prime},\nabla^{\prime})}$${(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})}$${(\bm{\breve{\Theta}^{\prime}},\bm{\breve{\mathcal{C}}}^{\prime})}$$\scriptstyle{F}$$\scriptstyle{F}$
This map $F$ is the formal Fourier transform. It acts independently on the Stokes circles: if we use the notation $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})=\sum_{I\in N}(I,\breve{\mathcal{C}}_{I})$ where $N$ is the set of active circles of $\bm{\breve{\Theta}}$ we have
$F\cdot\sum_{I\in N}(I,\breve{\mathcal{C}}_{I})=\sum_{I\in N}F\cdot(I,\breve{\mathcal{C}}_{I}).$
Furthermore, the irregular class $\bm{\breve{\Theta}}^{\prime}$ only depends on $\bm{\breve{\Theta}}$: there is a self-bijection of the set of all Stokes circles $\pi_{0}(\mathcal{I})$, that we will denote also by $F$, such that, if $\bm{\breve{\Theta}}=\sum n_{i}I_{i}$, then
$\bm{\breve{\Theta}}^{\prime}=F\cdot\bm{\breve{\Theta}}=\sum_{i}n_{i}(F\cdot I_{i}).$
This bijection is actually given by a Legendre transform. It can be computed explicitly, and takes slightly different forms depending on whether the Stokes circles are at finite distance or at infinity, and in the latter case on whether its slope is greater of less than 1. The next proposition describes the ramification orders, slopes and levels of the formal Fourier transform $F\cdot I$ of any Stokes circle $I$ as a function of those of $I$, as well as the transformation of the conjugacy classes of formal monodromies.
Proposition 2.14.
Let $I$ be a Stokes circle, and $\breve{\mathcal{C}}_{I}\in\operatorname{GL}_{n}(\mathbb{C})$ a conjugacy class. Let $\breve{\mathcal{C}}_{F\cdot I}\in\operatorname{GL}_{n}(\mathbb{C})$ be the conjugacy class such that $F\cdot(I,\breve{\mathcal{C}}_{I})=(F\cdot I,\breve{\mathcal{C}}_{F\cdot I})$. We have the following:
  1. (1)
    If $I$ i.e. of the form $\langle\alpha z\rangle_{\infty}$, with $\alpha\in\mathbb{C}$, then $F\cdot I=\langle 0\rangle_{\alpha}$, and $\breve{\mathcal{C}}_{F\cdot I}=-\breve{\mathcal{C}}_{I}$.
  2. (2)
    If $I$ is of slope $\leq 1$ at infinity, of the form $\langle\alpha z+q\rangle_{\infty}$, with $\alpha\in\mathbb{C}$, and $q\neq 0$ of slope $<1$ then $F\cdot I$ is of the form $F\cdot I=\langle\widetilde{q}\rangle_{\alpha}$, with $\operatorname{Irr}(\tilde{q})=s,\operatorname{ram}(\tilde{q})=r-s$, $\operatorname{Levels}(\tilde{q})=\frac{r}{r-s}\operatorname{Levels}(q)$. Also $\breve{\mathcal{C}}_{F\cdot I}=(-1)^{s}\breve{\mathcal{C}}_{I}$.
  3. (3)
    If $I$ is of slope >1 at infinity, with $\operatorname{ram}I=r$, $\operatorname{Irr}(I)=s$, then $F\cdot I$ satisifies $\operatorname{ram}(F\cdot I)=s-r$, $\operatorname{Irr}(F\cdot I)=s$, and $\operatorname{Levels}(F\cdot I)=\frac{r}{s-r}\operatorname{Levels}(I)$. Also $\breve{\mathcal{C}}_{F\cdot I}=(-1)^{s}\breve{\mathcal{C}}_{I}$.
  4. (4)
    If $I=\langle q\rangle_{a}$ is an irregular circle at finite distance for $a\in\mathbb{C}=\mathbb{P}^{1}\smallsetminus\{\infty\}$, with $q\neq 0$, $\operatorname{ram}(q)=r$, and $\operatorname{Irr}(q)=s$, then $F\cdot I$ is a circle of slope $\leq 1$, at infinity, of the form $\langle-az+\tilde{q}\rangle_{\infty}$, with $\operatorname{ram}(\tilde{q})=r+s,\operatorname{Irr}(\tilde{q})=s$, and $\operatorname{Levels}(\tilde{q})=\frac{r}{r+s}\operatorname{Levels}(q)$. Also $\breve{\mathcal{C}}_{F\cdot I}=(-1)^{s}\breve{\mathcal{C}}_{I}$.
  5. (5)
    If $I=\langle 0\rangle_{a}$ is a tame circle at finite distance, then $F\cdot I=\langle-az\rangle_{\infty}$, and $\breve{\mathcal{C}}_{F\cdot I}=-\breve{\mathcal{C}}_{I}$.
Proof.
All statements except those describing the sets of levels are already contained in the works [42, 37] establishing the stationary phase formula. For the levels, some of the statements are proven in [30] using results on singularities of curves. Let us discuss how to derive them from [18]. In the case where $I$ is at infinity and of slope $>1$, the statement for the levels can be deduced quite directly from Proposition 4.8 and the proof of Theorem 4.9 of [18]. Indeed, the proposition implies that implies that all possible exponents of $F\cdot I$ belong to a set determined by the levels of $I$, and that if $k$ is a level of $I$ then $kr/(r+s)$ is an exponent of $F\cdot I$ (because the coefficient of this exponent is nonzero), and the proof of Theorem 4.9 there establishes that these properties satisfied by the possible exponents of $F\cdot I$ imply that the exponents of $F\cdot I$ of the form $kr/(r+s)$ with $k\in\operatorname{Levels}(I)$ are the only ones where the sequence of common denominators of exponents of $F\cdot I$ strictly decreases, which precisely means that they are the levels of $F\cdot I$. The other cases will be completely similar, the only difference being the slope of $F\cdot I$. ∎
To fully determine the fission data of the Fourier transform, we must also describe the fission exponents between distinct Stokes circles of the Fourier transform which are at the same singularities. This can be explicitly done:
Proposition 2.15.
Let $I,J$ be two distinct circles. There are the following possibilities:
  • $\bullet$
    If $\pi(I)=\pi(J)=a\neq\infty$, then $\pi(F\cdot I)=\pi(F\cdot J)=\infty$ and
    • If $I$ and $J$ have no common part, then denoting $k=\operatorname{slope}(I)$, $l=\operatorname{slope}(k)$, we have $f_{I,J}=\max(\frac{k}{k+1},\frac{l}{l+1})$.
    • If $I$ and $J$ have a common part, denoting $k$ its slope, we have $f_{F\cdot I,F\cdot J}=\frac{1}{k+1}f_{I,J}$
  • $\bullet$
    If $\pi(I)\neq\infty,\pi(J)\neq\infty$ and $\pi(I)\neq\pi(J)$, then $\pi(F\cdot I)=\pi(F\cdot J)=\infty$ and $f_{F\cdot I,F\cdot J}=1$.
  • $\bullet$
    If $\pi(I)=\pi(J)=\infty$, $I,J$ are both of slope $\leq 1$ and are of the form $I=\langle\alpha z+q\rangle$, $J=\langle\alpha z+q^{\prime}\rangle$ with $\alpha\in\mathbb{C}$, and $q,q^{\prime}$ of slope $<1$, then $\pi(F\cdot I)=\pi(F\cdot J)=\alpha$ and
    • If $q$ and $q^{\prime}$ have no common part and respective slopes $k,l$, then $f_{F\cdot I,F\cdot J}=\max(\frac{k}{1-k},\frac{l}{1-l})$.
    • If $q$ and $q^{\prime}$ have a common part $q_{c}$, denoting $k$ its slope, we have $f_{F\cdot I,F\cdot J}=\frac{k}{1-k}f_{q,q^{\prime}}$.
  • $\bullet$
    If $\pi(I)=\pi(J)=\infty$, and $I,J$ are both of slope $>1$, then $\pi(F\cdot I)=\pi(F\cdot J)=\infty$ and
    • If $I$ and $J$ have no common part, then denoting $k=\operatorname{slope}(I)$, $l=\operatorname{slope}(J)$, we have $f_{I,J}=\max(\frac{k}{k-1},\frac{l}{l-1})$.
    • If $I$ and $J$ have a common part, let $k$ denote its slope, then $f_{F\cdot I,F\cdot J}=\frac{k}{k-1}f_{I,J}$
  • $\bullet$
    Otherwise, $\pi(I)\neq\pi(J)$.
Proof.
All statements where there is no common part involved follow directly from the previous proposition. For the statements in the cases where there is a common part, in the case where $I,J$ are at infinity and of slope $>1$ this is established in the proof of Theorem 5.9 of [18] (the word ‘fission exponent’ wasn’t used there: in terms of the terminology used there it corresponds to the highest number among the exponents of the leading term of the different parts of $F\cdot I$ and $F\cdot J$). The other cases will be similar. ∎
2.4.2. General symplectic $\operatorname{SL}_{2}(\mathbb{C})$ transformations
The Fourier-Laplace transform is part of a larger group of transformations acting on modules over the Weyl algebra. Indeed, to any matrix
$A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}$
in $\operatorname{SL}_{2}(\mathbb{C})$, we can associate an automorphism of the Weyl algebra $A_{1}=\mathbb{C}[z]\langle\partial_{z}\rangle$ given by $z\mapsto az+b\partial_{z},\partial_{z}\mapsto cz+d\partial_{z}$. This induces an action of the group $\operatorname{SL}_{2}(\mathbb{C})$ of symplectic transformations on modules over the Weyl algebra (see [37]).
The group of symplectic transformations is generated by three types of elementary transformations:
  • $\bullet$
    The Fourier-Laplace transform $F$, corresponding to the matrix $\begin{pmatrix}0&1\\ -1&0\end{pmatrix};$
  • $\bullet$
    Twists at infinity $T_{\lambda}$, for $\lambda\in\mathbb{C}$, corresponding to the matrix $\begin{pmatrix}1&\lambda\\ 0&1\end{pmatrix};$
  • $\bullet$
    Scalings $S_{\lambda}$, for $\lambda\in\mathbb{C}^{*}$, corresponding to the matrix $\begin{pmatrix}\lambda^{-1}&0\\ 0&\lambda\end{pmatrix}.$
The geometric interpretation of twists and scalings is the following. The twist $T_{\lambda}$ corresponds to taking the tensor product with the rank one module $(\mathbb{C}[z],\partial_{z}+\lambda z)$, and the scaling $S_{\lambda}$ corresponds to do the change of variable $z\mapsto z/\lambda$ on $\mathbb{P}^{1}$.
As for the Fourier transform, any element of $\operatorname{SL}_{2}(\mathbb{C})$ induces a transformation on modules over the Weyl algebra, and on irreducible connections on Zariski open subsets of the affine line. Moreover, there exists again a corresponding formal transformation (that we will also denote by $A$) on modified formal data with a similar structure. The action of twists and scalings on formal data and fission data is much simpler than for the Fourier transform:
Lemma 2.16.
Let $I$ be a Stokes circle, and $\breve{\mathcal{C}}_{I}\in\operatorname{GL}_{n}(\mathbb{C})$ a conjugacy class. Let $\breve{\mathcal{C}}_{T_{\lambda}\cdot I}\in\operatorname{GL}_{n}(\mathbb{C})$ be the conjugacy class such that $T_{\lambda}\cdot(I,\breve{\mathcal{C}}_{I})=(T_{\lambda}\cdot I,\breve{\mathcal{C}}_{T_{\lambda}\cdot I})$. We have the following:
  • $\bullet$
    If $\pi(I)=\infty$, then, $\pi(T_{\lambda}\cdot I)=\infty$, and if $I=\langle q\rangle_{\infty}$ we have $T_{\lambda}\cdot I=\langle q-\frac{\lambda}{2}z^{2}\rangle_{\infty}$. In particular $T_{\lambda}\cdot I$ has the same ramification order and set of levels as $I$, $\operatorname{slope}(T_{\lambda}\cdot I)=\max(\operatorname{slope}(\langle q\rangle),\operatorname{slope}(\langle q-\frac{\lambda}{2}z^{2}\rangle)$, for $\lambda\in\mathbb{C}$.
  • $\bullet$
    Otherwise if $\pi(I)\neq\infty$, $T_{\lambda}\cdot I=I$.
In any case $\breve{\mathcal{C}}_{T_{\lambda}\cdot I}=\breve{\mathcal{C}}_{I}$.
In brief, all exponential factors at infinity are shifted by $-\frac{\lambda}{2}z^{2}$ and all other exponential factors are left unchanged. As a consequence, the twist $T_{\lambda}$ preserves the global fission data and classes of global admissible deformations. The twist may change the slopes of the exponential factors at infinity, but leaves the slopes of Stokes circles at finite distance unchanged.
Lemma 2.17.
Let $I$ be a Stokes circle such that $\pi(i)=a\in\mathbb{P}^{1}$, and $\breve{\mathcal{C}}_{I}\in\operatorname{GL}_{n}(\mathbb{C})$ a conjugacy class. Let $\breve{\mathcal{C}}_{S_{\lambda}\cdot I}\in\operatorname{GL}_{n}(\mathbb{C})$ be the conjugacy class such that $S_{\lambda}\cdot(I,\breve{\mathcal{C}}_{I})=(S_{\lambda}\cdot I,\breve{\mathcal{C}}_{T_{\lambda}\cdot I})$. For $\lambda\in\mathbb{C}^{*}$,$\pi(S_{\lambda}\cdot I)=a/\lambda$, and writing $I=\langle q(z_{a}^{-1})\rangle$ with $q\in\mathbb{C}[z^{-1/r}]$, we have $S_{\lambda}\cdot I=\langle q(\lambda z_{a/\lambda}^{-1})\rangle$. Furthermore $\breve{\mathcal{C}}_{S_{\lambda}\cdot I}=\breve{\mathcal{C}}_{I}$.
As a consequence, for any $\lambda\in\mathbb{C}$, the scaling $S_{\lambda}$ preserves the global fission data and classes of global admissible deformations, as well as the slopes of all Stokes circles.
A crucial property of the diagram associated to a connection is that it is invariant under the $\operatorname{SL}_{2}(\mathbb{C})$ action.
Theorem 2.18 ([18]).
Let $(E,\nabla)$ be an irreducible connection on a Zariski open subset of $\mathbb{P}^{1}$ (different from a rank one connection with only a pole at infinity of order less than 2), and $A\in\operatorname{SL}_{2}(\mathbb{C})$. Then
$\Gamma(E,\nabla)=\Gamma(A\cdot(E,\nabla)).$
3. Nearby representations
In this section, we use the results of the previous paragraph to describe the formal data and classes of admissible deformations of the elements of the orbit of a connection $(E,\nabla)$ under the $\operatorname{SL}_{2}(\mathbb{C})$ action.
3.1. Generic form
Definition 3.1.
Let $I\in\pi_{0}(\mathcal{I})$ be a Stokes circle. We say that $I$ is of generic form if $\pi(I)=\infty$, and $\operatorname{slope}(q)\leq 2$.
If $I=\langle q\rangle_{\infty}$ is of generic form, the Fourier sphere coefficient $\lambda(I)$ of $I$ is the unique complex number $a\in\mathbb{C}$ such that $\langle\frac{a}{2}z^{2}+q\rangle$ has slope $<2$. Explicitly, this means that $q$ is of the form
$q=-\frac{a}{2}+q_{<2},$
where the exponential factor $q_{<2}$ has slope $<2$. We will denote $I_{<2}:=\langle q_{<2}\rangle$.
If $I$ is not of generic form, we set $\lambda(I)=\infty$.
For $A=\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{C})$ let $h_{A}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ denote the corresponding homography
$h_{A}:z\mapsto\frac{az+b}{cz+d}.$
We call the Fourier sphere, and denote by $\mathbf{P}^{1}$, the copy of $\mathbb{P}^{1}$ where the coefficients $\lambda(I)$ live (this terminology was introduced in [7]).
The following result says that $\operatorname{SL}_{2}(\mathbb{C})$ acts on the Fourier sphere by homographies:
Lemma 3.2.
Let $I\in\pi_{0}(\mathcal{I})$ be a Stokes circle, $a=\lambda(I)$ its Fourier sphere coefficient, and $A\in SL_{2}(\mathbb{C})$. Then
$\lambda(A\cdot I)=h_{A}(\lambda(I)).$
Proof.
In the case where $a\neq\infty$ and $h_{A}(a)\neq\infty$, this is the result of Proposition 4.13 of [18], and one checks that this is still true in the remaining cases. ∎
This immediately implies the following:
Corollary 3.3.
Let $I\in\pi_{0}(\mathcal{I})$ be a Stokes circle. Then for $A$ in an open dense subset of $\operatorname{SL}_{2}(\mathbb{C})$, $A\cdot I$ is of generic form.
Definition 3.4.
Let $\bm{\breve{\Theta}}$ be a global modified irregular class. We say that it is of generic form if each of its Stokes circles is of generic form. Similarly if $(E,\nabla)$ is an irreducible connection on a Zariski open subset of $\mathbb{P}^{1}$, we say that it is of generic form if its global modified irregular class is of generic form.
In particular a connection of generic form only has a singularity at infinity, i.e. $\bm{\breve{\Theta}}=\Theta_{\infty}$. A global modified irregular class $\bm{\breve{\Theta}}$ is of generic form if and only if $\lambda(\bm{\breve{\Theta}})\subset\mathbb{C}=\mathbb{P}^{1}\smallsetminus\{\infty\}$, where $\lambda(\bm{\breve{\Theta}})$ denotes the set of Fourier sphere coefficients of all its Stokes circles. This immediately implies:
Lemma 3.5.
Let $\bm{\breve{\Theta}}$ be an effective modified irregular class. Then for $A$ in an open dense subset of $\operatorname{SL}_{2}(\mathbb{C})$, $A\cdot\bm{\breve{\Theta}}$ is of generic form.
In particular, every $\operatorname{SL}_{2}(\mathbb{C})$ orbit contains an element of generic form. We can thus without loss of generality write any orbit as $\operatorname{SL}_{2}(\mathbb{C})\cdot(E,\nabla)$, with $(E,\nabla)$ of generic form.
As a preliminary step towards describing such orbits, let us first consider the case where $\Theta_{\infty}$ only has one Fourier sphere coefficient.
Proposition 3.6.
Let $(\Theta,\mathcal{C})$ be formal data at infinity with $\Theta$ of generic form, and such that $\lambda(\Theta)=\{a\}$ for some $a\in\mathbb{C}$. Let $A\in\operatorname{SL}_{2}(\mathbb{C})$. Let $\Theta^{<2}:=T_{-a}\cdot\Theta$ be the slope $<2$ irregular class obtained from $\Theta$ by subtracting $-\frac{a}{2}z^{2}$ to each of its Stokes circles. We have:
  • $\bullet$
    If $h_{A}(a)=\infty$, then $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $F\cdot(\Theta^{<2},\mathcal{C})$.
  • $\bullet$
    Otherwise, if $h_{A}(a)\neq\infty$, then $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $(\Theta,\mathcal{C})$ , or equivalently of $(\Theta^{<2},\mathcal{C})$.
Proof.
This is basically a consequence of the fact that any element in $\operatorname{SL}_{2}(\mathbb{C})$ can be factorised as a product of elementary transformations, and among those only the Fourier transform may act nontrivially on global fission data. More precisely, we can use lemma 4.15 of [18] to write down explicit factorisations of $A$. If $h_{A}(a)=\infty$, then we have $h_{AT_{a}F^{-1}}(\infty)=\infty$, so from this lemma, there exists $v\in\mathbb{C}^{*},\rho\in\mathbb{C}$ such that $AT_{a}F^{-1}=S_{v}T_{\rho}$, so that $A=S_{v}T_{\rho}FT_{-a}$. This implies $A\cdot(\Theta,\mathcal{C})=S_{v}T_{\rho}FT_{-a}\cdot(\Theta,\mathcal{C})=S_{v}T_{\rho}F\cdot(\Theta^{<2},\mathcal{C})$, and since twists and scalings do not change the equivalence class of admissible deformations $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $F\cdot(\Theta^{<2},\mathcal{C})$.
Let us now consider the case $h_{A}(a)\neq\infty$. If $h_{A}(\infty)\neq\infty$ then from the lemma $A$ admits a factorisation $A=S_{\nu}T_{\rho}$ with $v\in\mathbb{C}^{*},\rho\in\mathbb{C}$, which implies that $A\cdot\Theta$ is an admissible deformation of $\Theta$. Otherwise, from the lemma, $A$ admits a factorisation $A=T_{\mu}FS_{\nu}T_{\rho}$, with $v\in\mathbb{C}^{*},\mu,\rho\in\mathbb{C}$. In turn $\Theta^{\prime}:=S_{\nu}T_{\rho}\cdot\Theta$ is an admissible deformation of $\Theta$, and $\lambda(\Theta^{\prime})=\{a^{\prime}\}$, with $a^{\prime}=h_{S_{\nu}T_{\rho}}(a)$, and we have $a^{\prime}\neq 0$ otherwise we would have $h_{A}(a)=h_{S_{\nu}F}(a^{\prime})=\infty$. Now, applying the propositions 2.14 and 2.15 to $\Theta^{\prime}$, the fact that all exponential factors in $\Theta^{\prime}$ are of slope 2 and have the same nonzero leading coefficient implies that this is also the case for $F\cdot\Theta^{\prime}$, and that $F\cdot(\Theta^{\prime},\mathcal{C})$ is an admissible deformation of $(\Theta^{\prime},\mathcal{C})$. Since $T_{\mu}$ doesn’t change the equivalence class of admissible deformations, we finallly obtain that $A\cdot(\Theta,\mathcal{C})$ is an admissible deformation of $(\Theta,\mathcal{C})$, which concludes the proof. ∎
In particular, we can write down explicit formulas for the rank at infinity of the elements of the orbit of $\Theta$. Let $N$ be the set of Stokes circles of $\Theta$, and let us set:
$N^{+}:=\{\langle q\rangle\in N\;|\;\operatorname{slope}(q+\frac{a}{2}z^{2})>1\}\subset N,\qquad N^{-}:=\{\langle q\rangle\in N\;|\;\operatorname{slope}(q+\frac{a}{2}z^{2})\leq 1\}\subset N,$
so that $N=N^{+}\sqcup N^{-}$. Let us write the elements of $N^{-}$ as
$I^{-}_{j}=\langle-\frac{a}{2}z^{2}+b_{j}z+q^{-}_{j}\rangle$
for $j=1,\dots,|N^{-}|$, where $\operatorname{slope}(q^{-}_{j})<1$, and set $s^{-}_{j}:=\operatorname{Irr}(q^{-}_{j})$, $r^{-}_{j}:=\operatorname{ram}(q^{-}_{j})$. Similarly, we write
$I^{+}_{j}=\langle-\frac{a}{2}z^{2}+q^{+}_{j}\rangle,$
for $j=1,\dots,|N^{+}|$, with $\operatorname{slope}(q^{+}_{j})>1$, and set $s^{+}_{j}:=\operatorname{Irr}(q^{+}_{j})$, $r^{+}_{j}:=\operatorname{ram}(q^{+}_{j})$.
Lemma 3.7.
Let $n^{\pm}_{j}$ denote the multiplicities of the circles in $N^{\pm}$, and set $R^{+}_{j}:=s^{+}_{j}-r^{+}_{j}$, for $j=1,\dots,|N^{+}|$. Let $A\in\operatorname{SL}_{2}(\mathbb{C})$. We have
  • $\bullet$
[…]
  • $\bullet$
[…]
Proof.
This follows directly from the previous lemma and proposition 2.14, by observing that for the case $h_{A}(a)=\infty$, the Stokes circles in $N^{-}$ are sent by $A$ to finite distance (with the position of $A\cdot I_{j}^{-}$ determined by the coefficient $b_{j}$), while the circles in $N^{+}$ become circles of slope >1 at infinity, with $A\cdot I_{j}^{+}$ having ramification order $R_{j}^{+}$. ∎
3.2. The classes of nearby representations
We now discuss general orbits. Let $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ be global modified formal data, with $\bm{\breve{\Theta}}=\Theta_{\infty}$ be an effective irregular class of generic form. Let $a_{1},\dots a_{k}\in\mathbb{C}$ be the distinct Fourier sphere coefficients of its Stokes circles, i.e. we have $\lambda(\Theta_{\infty})=\{a_{1},\dots,a_{k}\}$.
Let $N$ be the set of Stokes circles of $\bm{\breve{\Theta}}$. For $I\in N$ we denote by $N_{I}\in\mathbb{N}_{>0}$ the multiplicity of $I$ in $\bm{\breve{\Theta}}$. We partition $N$ as follows: for $i=1,\dots,k$, we set
  • $\bullet$
    $N_{i}:=\{I\in N\;|\;\lambda(I)=a_{i}\}$,
  • $\bullet$
    $N_{i}^{+}:=\{\langle q\rangle\in N_{i}\;|\;\operatorname{slope}(q+\frac{a_{i}}{2}z^{2})>1\}\subset N_{i}$,
  • $\bullet$
    $N_{i}^{-}:=\{\langle q\rangle\in N_{i}\;|\;\operatorname{slope}(q+\frac{a_{i}}{2}z^{2})\leq 1\}\subset N_{i}$,
so that
$N=N_{1}\sqcup\dots\sqcup N_{k},\qquad N_{i}=N_{i}^{+}\sqcup N_{i}^{-}.$
We also set $N^{+}:=N_{1}^{+}\sqcup\dots\sqcup N_{k}^{+}$, and $N^{-}:=N_{1}^{-}\sqcup\dots\sqcup N_{k}^{-}$. For $i=1,\dots,k$, let $\Theta_{i}:=\sum_{I\in N_{i}}n_{I}I$ be the irregular class obtained from $\Theta_{\infty}$ obtained by discarding all circles with Fourier sphere coefficient different from $a_{i}$, $\mathcal{C}_{i}$ the corresponding collection of conjugacy classes, and let $\Theta^{<2}_{i}$ be the irregular class at infinity obtained from $\Theta_{i}$ by subtracting $-\frac{a_{i}}{2}z^{2}$ to each Stokes circle (in particular all circles of $\Theta_{i}$ have slope $<2$).
Theorem 3.8.
Let $A\in\operatorname{SL}_{2}(\mathbb{C})$, then
  • $\bullet$
    If $h_{A}(a_{i})=\infty$ for some $i\in\{1,\dots,k\}$, then $A\cdot(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ is an admissible deformation of
    $F\cdot(\Theta_{i}^{<2},\mathcal{C}_{i})+\sum_{l\neq i}(\Theta_{l},\mathcal{C}_{l}).$
  • $\bullet$
    Otherwise, if $h_{A}(a_{i})\neq\infty$ for each $i=1,\dots,k$, then $A\cdot(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ is an admissible deformation of $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$.
Proof.
This follows from applying proposition 3.6 to every $(\Theta_{i},\mathcal{C}_{i})$. ∎
This motivates the following definition:
Definition 3.9.
Let $\mathbf{T}$ be a fission tree of generic form, $N$ its set of leaves and $\bm{\mathcal{C}}$ the datum of a conjugacy class $\mathcal{C}_{i}\in\operatorname{GL}_{n_{i}}(\mathbb{C})$ for each leaf $i\in N$. Let $\Theta_{\infty}$ be an irregular class at infinity of generic form with fission tree $\mathbf{T}$. Keeping previous notations, for $i=1,\dots,k$, let $(\mathbf{F}_{i},\bm{\mathcal{C}}_{i})$ be the class of admissible deformations of the nonmodified formal data corresponding to the formal data
$F\cdot(\Theta_{i}^{<2},\mathcal{C})+\sum_{l\neq i}(\Theta_{l},\mathcal{C}).$
The definition makes sense, indeed it follows from Propositions 2.14 and 2.15 that the class of admissible deformations of $F\cdot(\Theta_{i}^{<2},\mathcal{C})+\sum_{l\neq i}(\Theta_{l},\mathcal{C})$ only depends on the fission tree $\mathbf{T}$. We call $(\mathbf{F}_{i},\bm{\mathcal{C}}_{i})$ the $i$-th nongeneric reading of $\mathcal{(}\mathbf{T},\bm{\mathcal{C}})$. The previous theorem can then be reformulated as follows:
Corollary 3.10.
Let $(\mathbf{\Theta},\bm{\mathcal{C}})$ be a wild Riemann surface with boundary data on $\mathbb{P}^{1}$ of generic form, with class of admissible deformations ($(\{\mathbf{T}\},\bm{\mathcal{C}})$. Then the classes of its nearby representations of $\mathcal{M}(\mathbf{\Theta},\bm{\mathcal{C}})$ are ($(\{\mathbf{T}\},\bm{\mathcal{C}})$ and $(\mathbf{F}_{i},\bm{\mathcal{C}}_{i})$ for $i=1,\dots k$.
By abuse of language, we will sometimes allow ourselves to say that $\{\mathbf{T}\}$ and $\mathbf{F}_{i}$, for $i=1,\dots,k$ are the nearby representations of $\mathbf{T}$.
Finally, all this allows us to have a well-defined notion of enriched tree associated to any irregular connection on $\mathbb{P}^{1}$.
Proposition 3.11.
Let $(E,\nabla)$ be an arbitrary algebraic connection on a Zariski open subset of $\mathbb{P}^{1}$, with global modified irregular class $(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$, $A\in\operatorname{SL}_{2}(\mathbb{C})$ such that $A\cdot(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$ is of generic form. Let $\mathbf{T}$ be the (short) fission tree of $A\cdot\bm{\breve{\Theta}}$, and $\bm{\mathcal{C}}$ the collection of conjugacy classes of $A\cdot(\bm{\breve{\Theta}},\bm{\breve{\mathcal{C}}})$. Then the pair $(\mathbf{T},\bm{\mathcal{C}})$ does not depend on the choice of $A$.
Proof.
This is an immediate consequence of Thm. 3.8. ∎
Definition 3.12.
Keeping the setup of the proposition, we say that $(\mathbf{,}\bm{\mathcal{C}})$ is the enriched tree of $(E,\nabla)$ and write $\mathscr{T}(E,\nabla):=(\mathbf{T},\bm{\mathcal{C}})$.
Prop. 3.11 implies that the enriched tree is invariant under symplectic transformations.
Corollary 3.13.
If $(E,\nabla)$ is irreducible and is not a rank one connection with just one pole of order less than 2 at infinity, so that $A\cdot(E,\nabla)$ is well-defined for any $A\in\operatorname{SL}_{2}(\mathbb{C})$, we have
$\mathscr{T}(A\cdot(E,\nabla))=\mathscr{T}(E,\nabla).$
3.3. Properties of classes of nearby representations
We can provide explicit formulas for the rank of the nearby representations. For $i=1,\dots,k$, let us denote by $\langle q^{+}_{i,j}\rangle$ the elements of $N_{i}^{+}$, and write $\langle q^{+}_{i,j}\rangle=\langle-\frac{a_{i}}{2}z^{2}+q^{\prime+}_{ij}\rangle$, with $q^{\prime+}_{ij}$ having ramification $r^{+}_{ij}$ and slope $s^{+}_{ij}/r^{+}_{ij}>1$. Similarly, let us denote by $\langle q^{-}_{i,j}\rangle$ the elements of $N_{i}^{-}$, and write $\langle q^{-}_{i,j}\rangle=\langle-\frac{a_{i}}{2}z^{2}+b_{ij}z+q^{\prime-}_{ij}\rangle$, with $q^{\prime-}_{ij}$ having ramification $r^{+}_{ij}$ and slope $s^{+}_{ij}/r^{+}_{ij}<1$. Let $n_{ij}^{\pm}$ be the multiplicities of the active circles.
Proposition 3.14.
Let us set $R^{+}_{ij}=s^{+}_{ij}-r^{+}_{ij}$.
  1. (1)
    The generic representation has rank
    $r_{gen}=\sum_{i}\left(\sum_{j}n^{+}_{ij}r^{+}_{ij}+\sum_{j}n^{-}_{ij}r^{-}_{ij}\right).$
  2. (2)
    The $i$-th nongeneric representation has rank
    $r_{i}=\sum_{j}n^{+}_{ij}R^{+}_{ij}+\sum_{l\neq i}\left(\sum_{j}n^{+}_{lj}r^{+}_{lj}+\sum_{j}n^{-}_{lj}r^{-}_{lj}\right).$
Proof.
This follows directly from Lemma 3.7. ∎
We can also express the number of singularities at finite distance of the nearby representations:
Proposition 3.15.
Let $(E,\nabla)$ of generic form, and $(E^{\prime},\nabla^{\prime})\in\operatorname{SL}_{2}(\mathbb{C})\cdot(E,\nabla)$. If $(E^{\prime},\nabla^{\prime})$ is in the $i$-th nongeneric class of representations, its number of singularities at finite distance is the number of distinct coefficients $b_{ij}$ among the circles in $N_{i}^{-}$.
Proof.
This is a consequence of item (2) of Prop. 2.14: the position of the singularity above which $F\cdot\langle q^{-}_{i,j}\rangle$ lies is determined by the coefficient $b_{ij}$. ∎
In particular, if $N_{i}^{-}=\emptyset$, then the $i$-th nongeneric class of representations only has a singularity at infinity. In general, the class of nearby representations all have different ranks and number of singularities.
Finally, we can also read from the generic tree when two nongeneric representations have the same forest.
Proposition 3.16.
Keeping previous notations, for $i,j=1,\dots k$, with $i\neq j$, the fission forests $\mathbf{F}_{i}$ and $\mathbf{F}_{j}$ are isomorphic if and only if the $i$-th and $j$-th principal subtrees of the fission tree $\mathbf{T}$ are isomorphic.
Proof.
Since isomorphism classes of fission trees characterise admissible deformations, it is equivalent to show that the $i$-th and the $j$-th nongeneric forests are isomorphic if and only if $\Theta_{i}^{<2}$ and $\Theta_{j}^{<2}$ are admissible deformations of each other. Let us first assume that $\Theta_{i}^{<2}$ and $\Theta_{j}^{<2}$ are admissible deformations of each other. We have to be careful that in general if two irregular class $\Theta$ and $\Theta^{\prime}$ at infinity are admissible deformations of each other, it is not always true that so are their Fourier transform, because if $\langle q\rangle$ is a Stokes circle for $\Theta$ and $\langle q^{\prime}\rangle$ the Stokes circle for $\Theta^{\prime}$ corresponding to $\langle q\rangle$ under some admissible deformation, their slope may differ which may imply that $F\cdot\langle q\rangle$ and $F\cdot\langle q^{\prime}\rangle$ are not admissible deformations of each other. However here because all Stokes circles of $\Theta_{i}^{<2}$ and $\Theta_{j}^{<2}$ are of slope $<2$ this cannot happen and we have that $F\cdot\Theta_{i}^{<2}$ and $F\cdot\Theta_{j}^{<2}$ are admissible deformations of each other. In turn, the irregular classes $\Theta_{i}^{<2}+\sum_{l\neq i}\Theta_{l}$ and $\Theta_{j}^{<2}+\sum_{l\neq j}\Theta_{l}$ are admissible deformations of each other, so the $i$-th and $j$-th representations coincide. Conversely, if the $i$-th and $j$-th classes of representations coincide, then in particular $\sum_{l\neq i}\Theta_{l}$ and $\sum_{l\neq j}\Theta_{l}$ are admissible deformations of each other, which implies that $\Theta_{i}$ and $\Theta_{j}$ are admissible deformations of each other, and the conclusion follows. ∎
Corollary 3.17.
Let $\bm{\breve{\Theta}}$ be an effective modified global irregular class. Then the number of isomorphism classes of fission forests among the forests of elements of the orbit $\operatorname{SL}_{2}(\mathbb{C})\cdot\bm{\breve{\Theta}}$ is equal to 1 plus the number of distinct isomorphism classes of fission trees among the principal subtrees of its generic fission tree.
Remark 3.1.
The classes of nearby representations arise from the action of $\operatorname{SL}_{2}(\mathbb{C})$ on connections on the affine line. However, as already mentioned in the introduction, there is a larger class of natural operations on connections: indeed one can consider Möbius transformations, which have the effect of changing the point at infinity in the identification $\mathbb{P}^{1}=\mathbb{C}\cup\{\infty\}$, as well as the operation consisting in tensoring a connection with an arbitrary rank one local system on a Zariski open subset of $\mathbb{P}^{1}$. This larger class of operations on connections on $\mathbb{P}^{1}$ plays an essential role in the extension due to Deligne-Arinkin [17, 3] of the Katz algorithm for rigid local systems to the case of irregular connections, which provides a way to reduce any irreducible rigid connection on $\mathbb{P}^{1}$ to the trivial rank one connection by repeated application of one of these operations. Since the symplectic $\operatorname{SL}_{2}(\mathbb{C})$ transformations don’t commute with these other operations in general, the latter do not preserve the diagram or the classes of nearby representations.
4. Basic representations for Painlevé moduli spaces
In this section, as a set of examples illustrating our general discussion, we describe the basic representations of all Painlevé moduli spaces, obtained from the standard rank 2 Lax representation of each Painlevé equation. Recall that, while there were not first discovered in this way historically, all Painlevé equations can be obtained from isomonodromic deformations. Each Painlevé equation arises from isomonodromic deformations of connections with a specific type of singularity data, i.e. different Painlevé equations correspond to different nonabelian Hodge spaces. In the standard way to derive the Painlevé equations from isomonodromy [33], all Lax representations are of rank 2, and in all cases except for Painlevé VI, these Lax representations have at least one irregular singularity.
For each case, starting from this standard Lax representation (in its abstract form, i.e. consisting in the data of a wild Riemann surface), we will draw the corresponding generic fission tree of minimal rank, describe the classes of nearby representations, and discuss when some nongeneric representations have the same forest. While the cases of Painlevé VI, V, IV all fall in the simply-laced case and were discussed by Boalch in [7], it is still worth discussing them briefly from the point of view of generic fission trees. On the other hand, describing all basic representations for Painlevé III, II, I as well as for the moduli spaces corresponding to degenerate versions of Painlevé III really requires using our general framework. Doing this, we recover among the classes of nearby representations many known alternative Lax representations for various Painlevé equations, e.g. from [25, 34, 35] and the other representations can be viewed as new abstract Lax representations.
Notice that there is the following subtlety: since taking the Fourier transform doesn’t always commute with doing admissible deformations, different connections having the same fission forest in general do not have the same classes of nearby representations, nor the same diagram. In the cases where this happens, among all possible wild Riemann surfaces with boundary data having the fission forest of the standard Painlevé Lax representation, we choose one such that its generic representation has minimal rank. Making this choice is necessary to allow us to recover several known alternative Lax representations for Painlevé equations and to recover the diagrams for Painlevé equations obtained in [14], all equal or related to affine Dynkin diagrams. Other choices would lead to different diagrams, and for each choice all nearby representations could also be explicitly determined, yielding even more abstract Lax representations, typically of higher rank.
To draw the fission forests and the generic fission tree, we follow the conventions of [11]. The root of the trees is indicated by a black square vertex and the heights of vertices of the trees are indicated to the left of the figure. In [11] it was necessary to introduce different types of vertices, represented by different colors (black or white): we haven’t recalled the corresponding definitions here since this distinction doesn’t play a role for our present purposes, but will still represent the vertices with their right colours for the sake of consistency. The multiplicity of the leaves (i.e. the rank the corresponding conjugacy classes) is equal to 1 unless indicated otherwise.
To draw the diagrams, we follow the same conventions as in [18]: unless otherwise specified the multiplicity of each vertex is 1, negative edges or loops are drawn in dashed lines, and for multiple edges/loops the multiplicity in indicated at their middle.
The generic fission trees as well as the rank and number of singularities of each class of basic representations for each Painlevé moduli space are listed below. In the next paragraphs, we explain in more detail how to obtain these classes of representations starting from the standard Lax representations.
Painlevé I
22225/3535/35 / $34/$ 311112/3232/32 / $31/$ 3
Rank Number of singularities
Generic rep. 3 1
Nongeneric rep. 2 1
Painlevé II
111122223/2323/23 / 2
Rank Number of singularities
Generic rep. 3 1
1st nongeneric rep. 2 1
2nd nongeneric rep. 2 2
Painlevé III
111122221/2121/21 / 2
Rank Number of singularities
Generic rep. 4 1
1st nongeneric rep. 2 3
2st nongeneric rep. 2 2
Degenerate Painlevé III
111122221/2121/21 / 2
Rank Number of singularities
Generic rep. 4 1
Nongeneric reps. 2 2
Doubly degenerate Painlevé III
111122221/2121/21 / 21/3131/31 / 3
Rank Number of singularities
Generic rep. 5 1
1st nongeneric rep. 2 2
2st nongeneric rep. 3 2
Painlevé IV
22221111
Rank Number of singularities
Generic rep. 3 1
Nongeneric reps. 2 2
Painlevé V
11112222
Rank Number of singularities
Generic rep. 4 1
Nongeneric reps. 2 3
Painlevé VI
111122222222
Rank Number of singularities
Generic rep. 5 1
1st nongeneric rep. 2 4
2st nongeneric rep. 3 2
4.1. Painlevé I
The standard rank 2 Painlevé I Lax representation has a single irregular singularity, at infinity, given by a pole of order 4, but such that the leading term is nilpotent, which leads to a Turritin-Levelt normal form having a single twisted Stokes circle of slope $5/2$.
The corresponding fission forest (having a single tree) is
5/2525/25 / 222223/2323/23 / $21/$ 2
It admits representatives given by irregular classes of the form
$\bm{\breve{\Theta}}^{I}_{std}=\langle\alpha z^{5/2}\rangle_{\infty},$
for any $\alpha\neq 0$.
Let us study the classes of representations corresponding to the orbit of $\bm{\breve{\Theta}}_{std}^{I}$. The generic class of representations has rank 3, and admits representatives of the form
$\bm{\breve{\Theta}}^{I}_{gen}=\langle\lambda z^{2}+\beta z^{5/3}\rangle_{\infty},$
with $\beta\neq 0$.
The generic tree is
22225/3535/35 / $34/$ 311112/3232/32 / $31/$ 3
There is a single Fourier sphere coefficient here so $N=N_{1}=N_{1}^{+}$, and the standard representation corresponds to the only nongeneric class of representations. Notice that the generic class of representations corresponds to the (abstract version of the more explicit) alternative Lax representation for Painlevé I found in [35].
The corresponding diagram only has one vertex and one loop, and is drawn below:
The two classes of representations correspond to different readings of the diagram.
4.2. Painlevé II
The standard rank 2 Painlevé II Lax representation has only one irregular singularity, at infinity, given by a pole of order 4.
The corresponding fission forest is
111122223333
It admits representatives given by irregular classes of the form:
$\bm{\Theta}^{II}=\bm{\breve{\Theta}}^{II}=\langle\alpha z^{3}\rangle_{\infty}+\langle\beta z^{3}\rangle_{\infty},$
with $\alpha\neq\beta$.
To have a minimal rank of the generic class of representations and recover the other known Lax representations for Painlevé II, we need to assume that $\beta=0$, to which we can always reduce up to tensoring with a rank one local system (but beware this twist is by an exponential factor of slope 3, so does not belong to the $\operatorname{SL}_{2}(\mathbb{C})$ action. However, since it is at infinity, it preserves the diagram). Let us thus consider an irregular class of the form
$\bm{\breve{\Theta}}^{II}_{std}=\langle\alpha z^{3}\rangle_{\infty}+\langle 0\rangle_{\infty},$
with $\alpha\neq 0$ and study the classes of its nearby representations.
The generic class of representations has rank 3, and admits representatives of the form
$\bm{\breve{\Theta}}^{II}_{gen}=\langle\lambda z^{2}+\alpha z^{3/2}\rangle_{\infty}+\langle\mu z^{2}\rangle_{\infty},$
with $\lambda\neq\mu$, $\alpha\neq 0$.
The generic fission tree is
111122223/2323/23 / 2
We have $k=2$ Fourier sphere coefficients, $N_{1}=N_{1}^{+}=\{\langle\lambda z^{2}+\alpha z^{3/2}\rangle_{\infty}\}$, $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}\rangle_{\infty}\}$. The two principal subtrees are not isomorphic, so the two nongeneric classes of representations have distinct forests.
The standard representation corresponds to the nongeneric class of representations for $i=1$. The second nongeneric class of representations also has rank 2, and admits representatives of the form
$\bm{\breve{\Theta}}^{II}_{2}=\langle\alpha z^{3/2}\rangle_{\infty}+\langle 0\rangle_{0}.$
We notice that it corresponds precisely the other known Lax representation for Painlevé II due to Flaschka-Newell [25], also sometimes referred to as degenerate Painlevé IV.
Finally the diagram is
and the basic representations correspond to different readings of the diagram, as indicated on the figure below:
$\langle\alpha z^{3}\rangle_{\infty}$$\langle 0\rangle_{\infty}$$\bm{\breve{\Theta}}^{II}_{std}$$\langle\lambda z^{2}+\alpha z^{3/2}\rangle_{\infty}$$\langle\mu z^{2}\rangle_{\infty}$$\bm{\breve{\Theta}}^{II}_{gen}$$\langle\alpha z^{3/2}\rangle_{\infty}$$\langle 0\rangle_{0}$$\bm{\breve{\Theta}}^{II}_{2}$
4.3. Painlevé III
The standard rank 2 Painlevé III Lax representation has two irregular singularities, both corresponding to a second order pole, one of them being at infinity.
The fission forest is
$1$
The general form of a global irregular class with this fission forest and a singularity at infinity is  [^1] We will denote Painlevé III, degenerate Painlevé III and doubly degenerate Painlevé III by $III(2)$, $III(1)$ and $III(0)$ respectively, the numbers 2,1,0 referring to the number of parameters of each equation.
$\bm{\Theta}^{III(2)}_{std}=\langle\alpha z\rangle_{\infty}+\langle\beta z\rangle_{\infty}+\langle\alpha^{\prime}z^{-1}_{a}\rangle_{a}+\langle\beta^{\prime}z_{a}^{-1}\rangle_{a},$
with $\alpha,\beta,\alpha^{\prime},\beta^{\prime}\in\mathbb{C}$, $a\neq\infty$, $\alpha\neq\beta$, $\alpha^{\prime}\neq\beta^{\prime}$.
If $(E,\nabla)$ is a connection of this type, we can always apply a twist by a rank one local system to set $\beta^{\prime}$ to zero, to obtain a connection with a modified global irregular class of the form (whose fission forest is the one drawn on Fig. LABEL:fig:PIII_tree_and_diagram)
$\bm{\breve{\Theta}}^{III(2)}_{std}=\langle\alpha z\rangle_{\infty}+\langle\beta z\rangle_{\infty}+\langle\alpha^{\prime}z_{a}^{-1}\rangle_{a}.$
with $\alpha^{\prime}\neq 0$.
Let us study the classes of nearby representations of $\bm{\breve{\Theta}}^{III(2)}_{std}$. The generic class of representations has rank 4, and admits representatives of the following form
$\bm{\breve{\Theta}}^{III(2)}_{gen}=\langle\lambda z^{2}+\alpha z\rangle_{\infty}+\langle\lambda z^{2}+\beta z\rangle_{\infty}+\langle\mu z^{2}+\gamma z^{1/2}\rangle_{\infty},$
with $\lambda\neq\mu$, $\alpha\neq\beta$, $\gamma\neq 0$.
We have $k=2$ Fourier sphere coefficients, $N_{1}=N_{1}^{-}=\{\langle\lambda z^{2}+\alpha z\rangle_{\infty},\langle\lambda z^{2}+\beta z\rangle_{\infty}\}$, $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}+\gamma z^{1/2}\rangle_{\infty}\}$.
The corresponding fission tree is
111122221/2121/21 / 2
Since the two principal subtrees are not isomorphic, the forests of the two corresponding nongeneric representations do not coincide. The standard representation is the nongeneric representation for $i=1$. The other nongeneric class of representations has rank 2 and corresponds to irregular classes of the form
$\bm{\breve{\Theta}}^{III(2)}_{2}=\langle\alpha z^{1/2}\rangle_{\infty}+\langle 0\rangle_{a}+\langle 0\rangle_{b},$
with $a\neq b$, $\alpha\neq 0$. It is none other but the Flaschka-Newell Lax representation for Painlevé III [25], also sometimes referred to as degenerate Painlevé V.
Finally, the diagram is
and the classes of basic representations correspond to different readings, as represented on the figure below
$\langle\lambda z^{2}+\alpha z\rangle_{\infty}$$\langle\mu z^{2}+\gamma z^{1/2}\rangle_{\infty}$$\langle\lambda z^{2}+\beta z\rangle_{\infty}$$\bm{\breve{\Theta}}^{III(2)}_{gen}$$\langle\alpha z\rangle_{\infty}$$\langle\alpha^{\prime}z_{a}\rangle_{a}$$\langle\beta z\rangle_{\infty}$$\bm{\breve{\Theta}}^{III(2)}_{std}$$\langle 0\rangle_{a}$$\langle\alpha z^{1/2}\rangle_{\infty}$$\langle 0\rangle_{b}$$\bm{\breve{\Theta}}^{III(2)}_{2}$
4.4. Degenerate Painlevé III
The standard Lax representation for degenerate Painlevé III, as for nondegenerate Painlevé III, has two irregular singularities, one of them at infinity, both corresponding to poles of order 2, but with a nilpotent leading term for the singularity at infinity, giving rise in the Turritin-Levelt normal form to a single Stokes circle of slope $1/2$ [40].
The general form of a global irregular class of this type is
$\bm{\Theta}^{III(1)}_{std}=\langle\beta z^{1/2}\rangle_{\infty}+\langle\alpha z_{a}^{-1}\rangle_{a}+\langle\alpha^{\prime}z_{a}^{-1}\rangle_{a},$
with $a\neq\infty$, $\beta\neq 0$, and $\alpha\neq\alpha^{\prime}$.
The corresponding fission forest is
11111/2121/21 / 2
Given a connection of this type, we can always apply a twist by a rank one local system to set $\alpha^{\prime}=0$ and kill the formal monodromy of $\langle\alpha^{\prime}z_{a}^{-1}\rangle_{a}$, and obtain a connection with a global modified irregular class of the form:
$\bm{\breve{\Theta}}^{III(1)}_{std}=\langle\beta z^{1/2}\rangle_{\infty}+\langle\alpha z^{-1}\rangle_{a},$
with $\beta\neq 0$, $\alpha\neq 0$.
Now let us consider the classes of nearby representations of $\bm{\breve{\Theta}}^{III(1)}_{std}$. The generic class of representations has rank 4, and admits representatives of the form
$\bm{\breve{\Theta}}^{III(1)}_{gen}=\langle\lambda z^{2}+\alpha z^{1/2}\rangle_{\infty}+\langle\mu z^{2}+\beta z^{1/2}\rangle_{\infty}$
with $\lambda\neq\mu$, $\alpha\neq 0$ and $\beta\neq 0$.
The generic fission tree is
111122221/2121/21 / 2
We have $k=2$ Fourier sphere coefficients, $N_{1}=N_{1}^{-}=\{\langle\lambda z^{2}+\alpha z^{1/2}\rangle_{\infty}\}$, and $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}+\beta z^{1/2}\rangle_{\infty}\}$.
Furthermore, the two principal subtrees are isomorphic, so the two nongeneric classes of representations have the same forest, and there are only two distinct forests, the generic one and the nongeneric one.
Finally, the diagram is
444411-1- $-1$
and the distinct forests correspond to different types of readings of the diagram, as indicated on the figure below:
[…]
4.5. Doubly degenerate Painlevé III
The Lax representation for the doubly degenerate Painlevé equation has two irregular singularities, one of them at infinity, both corresponding to a second order pole, but with a nilpotent leading term at each pole so that at each pole, in the Turritin-Levelt decomposition, there is a single twisted Stokes circle of slope 1/2 [40]. This corresponds to a global (modified or not) irregular class of the form
$\bm{\Theta}^{III(0)}_{std}=\bm{\breve{\Theta}}^{III(0)}_{std}=\langle\alpha z^{1/2}\rangle_{\infty}+\langle\beta z^{-1/2}\rangle_{a},$
with $a\neq\infty$, $\alpha\neq 0$, $\beta\neq 0$.
The corresponding forest is
$1/2$
Let us study the orbit $\operatorname{SL}_{2}(\mathbb{C})\cdot\bm{\breve{\Theta}}^{III(0)}_{std}$. The generic representation has rank 5 and admits representatives of the form
$\bm{\breve{\Theta}}^{III(0)}_{gen}=\langle\lambda z^{2}+\alpha z^{1/3}\rangle_{\infty}+\langle\mu z^{2}+\beta z^{1/2}\rangle_{\infty},$
with $\lambda\neq\mu$, $\alpha\neq 0$, $\beta\neq 0$.
The generic tree is
111122221/2121/21 / 21/3131/31 / 3
We have $k=2$ Fourier sphere coefficients, $N_{1}=N_{1}^{-}=\{\langle\lambda z^{2}+\alpha z^{1/3}\rangle_{\infty}\}$ and $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}+\beta z^{1/2}\rangle_{\infty}\}$. Since the principal subtrees are not isomorphic, the two nongeneric forests are distinct. Besides the class of the standard representation, corresponding to $i=1$, the other nongeneric one has rank 3 and admits representatives of the form
$\bm{\breve{\Theta}}^{III(0)}_{2}=\langle\alpha z^{1/3}\rangle_{\infty}+\langle\beta z^{-1}\rangle_{0},$
with $\alpha\neq 0$, $\beta\neq 0$.
Finally the diagram is
666611-1- 133-3- 3
and the three classes of representations correspond to different readings of the diagram, as indicated on the figure below:
[…]
4.6. Painlevé IV
The standard rank 2 Painlevé IV Lax representation has an irregular singularity at infinity corresponding to a pole of order 3, and a regular singularity at finite distance.
The corresponding fission forest is
11112222
Any connection of this type can be brought, by applying a twist to kill a part of the formal monodromy at the regular singularity, to a connection with a modified irregular class of the form
$\bm{\breve{\Theta}}^{IV}_{std}=\langle\alpha z^{2}\rangle_{\infty}+\langle\beta z^{2}\rangle_{\infty}+\langle 0\rangle_{a},$
with $\alpha\neq\beta$.
Let us describe the classes of representations of $\bm{\breve{\Theta}}^{IV}_{std}$. The generic representation has rank 3. It admits representatives of the form
$\bm{\breve{\Theta}}^{IV}_{gen}=\langle\alpha z^{2}\rangle_{\infty}+\langle\beta z^{2}\rangle_{\infty}+\langle\gamma z^{2}\rangle_{\infty},$
with $\alpha,\beta,\gamma$ pairwise distinct. It coincides with the rank 3 Lax representation found in [34].
The fission tree of the generic representation is
22221111
We have $k=3$ Fourier sphere coefficients, $N_{1}=N_{1}^{-}=\{\langle\alpha z^{2}\rangle_{\infty}\}$, $N_{2}=N_{2}^{-}=\{\langle\beta z^{2}\rangle_{\infty}\}$, and $N_{3}=N_{3}^{-}=\{\langle\gamma z^{2}\rangle_{\infty}\}$.
We observe that the three principal subtrees of the generic fission forest are isomorphic. This implies that the forests of the three nongeneric representations coincide (this reflects the fact that $\alpha,\beta,\gamma$ have entirely symmetric roles), so there are only two distinct forests, the generic one and a nongeneric one.
Finally, the diagram is the following:
and the two forests correspond to different types of readings of the diagram, as indicated on the figure below:
$\langle\alpha z^{2}\rangle_{\infty}$$\langle\beta z^{2}\rangle_{\infty}$$\langle\gamma z^{2}\rangle_{\infty}$$\bm{\breve{\Theta}}^{IV}_{gen}$$\langle\alpha z^{2}\rangle_{\infty}$$\langle\beta z^{2}\rangle_{\infty}$$\langle 0\rangle_{a}$$\bm{\breve{\Theta}}^{IV}_{std}$
4.7. Painlevé V
The standard Painlevé V Lax representation has rank 2, one irregular singularity at infinity corresponding to a pole of order 2, and two regular singularities at finite distance.
The corresponding fission forest is
$1$
If $(E,\nabla)$ is a connection of this type, we can always apply a twist to kill a part of the formal monodromy at the regular singularities, to obtain a connection with a modified global irregular class of the form:
$\bm{\breve{\Theta}}^{V}_{std}=\langle\alpha z\rangle_{\infty}+\langle\beta z\rangle_{\infty}+\langle 0\rangle_{a}+\langle 0\rangle_{b},$
with $\alpha\neq\beta$, $a\neq b$.
Let us describe the classes of representations of $\bm{\breve{\Theta}}^{V}_{std}$. The generic representation has rank 4 and corresponds to irregular classes of the form
$\bm{\breve{\Theta}}^{V}_{gen}=\langle\lambda z^{2}+\alpha z\rangle_{\infty}+\langle\lambda z^{2}+\beta z\rangle_{\infty}+\langle\mu z^{2}+az\rangle_{\infty}+\langle\mu z^{2}+bz\rangle_{\infty}$
with $\lambda\neq\mu$, $\alpha\neq\beta$, $a\neq b$.
The generic fission tree is thus
11112222
We have $k=2$ Fourier sphere coefficients, $N_{1}=N_{1}^{-}=\{\langle\lambda z^{2}+\alpha z\rangle_{\infty},\langle\lambda z^{2}+\beta z\rangle_{\infty}\}$ and $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}+az\rangle_{\infty},\langle\mu z^{2}+bz\rangle_{\infty}\}$. Furthermore, the two principal subtrees corresponding to $N_{1}$ and $N_{2}$ are isomorphic, so the two nongeneric classes of representations have in fact the same forest. In turn, there are only two distinct forests of basic representations, the generic forest and the nongeneric one.
The diagram associated to $\bm{\breve{\Theta}}^{V}_{std}$ is
and the two forests correspond to different types of readings of the diagram, as indicated on the figure below:
$\langle\lambda z^{2}+\alpha z\rangle_{\infty}$$\langle\lambda z^{2}+\beta z\rangle_{\infty}$$\langle\mu z^{2}+az\rangle_{\infty}$$\langle\mu z^{2}+bz\rangle_{\infty}$$\bm{\breve{\Theta}}^{V}_{gen}$$\langle\alpha z\rangle_{\infty}$$\langle\beta z\rangle_{\infty}$$\langle 0\rangle_{a}$$\langle 0\rangle_{b}$$\bm{\breve{\Theta}}^{V}_{std}$
4.8. Painlevé VI
The standard rank 2 Lax representation for Painlevé VI has 4 regular singularities, one of them at infinity. It corresponds to a global irregular class of the form
$\bm{\Theta}^{VI}_{std}=2\langle 0\rangle_{\infty}+2\langle 0\rangle_{a}+2\langle 0\rangle_{b}+2\langle 0\rangle_{c}.$
with poles at finite distance $a,b,c,$ pairwise distinct.
The fission forest is
Furthermore, the formal monodromy at every singularity has two distinct eigenvalues. Up to tensoring with a rank one local system with regular singularities at $a,b,c$, we may assume that one of these eigenvalues is equal to 1. The corresponding modified irregular class is then
$\bm{\breve{\Theta}}^{VI}_{std}=2\langle 0\rangle_{\infty}+\langle 0\rangle_{a}+\langle 0\rangle_{b}+\langle 0\rangle_{c}.$
Let us describe the classes of nearby representations of $\bm{\breve{\Theta}}^{VI}_{std}$. The generic class of representations can be reached from $\bm{\breve{\Theta}}^{VI}_{std}$ by applying a twist followed by the Fourier transform. It has rank 5 and corresponds to irregular classes of the form:
$\bm{\breve{\Theta}}^{VI}_{gen}=2\langle\lambda z^{2}\rangle_{\infty}+\langle\mu z^{2}+\alpha z\rangle_{\infty}+\langle\mu z^{2}+\beta z\rangle_{\infty}+\langle\mu z^{2}+\gamma z\rangle_{\infty},$
with $\lambda\neq\mu$, $\alpha,\beta,\gamma$ pairwise distinct.
The corresponding fission tree is
111122222222
The number of Fourier sphere coefficients is $k=2$, and we have $N_{1}=N_{1}^{-}=\{\langle\lambda z^{2}\rangle\}$, $N_{2}=N_{2}^{-}=\{\langle\mu z^{2}+\beta z\rangle,\langle\mu z^{2}+\gamma z\rangle\}$.
In turn there are three distinct classes of representations. The standard representation corresponds to the nongeneric class of representations for $i=1$. The other nongeneric class of representations, for $i=2$, has rank 3 and admits representatives of the form
$\bm{\breve{\Theta}}^{VI}_{2}=2\langle 0\rangle_{0}+\langle\alpha z\rangle_{\infty}+\langle\beta z\rangle_{\infty}+\langle\gamma z\rangle_{\infty},$
with $\alpha,\beta,\gamma$ pairwise distinct. It corresponds to the Harnad dual [28] of the standard Lax representation.
The diagram associated to $\bm{\breve{\Theta}}^{VI}_{std}$ is the affine D4 Dynkin diagram drawn below
where the right node is not coloured to indicate that it does not belong to the core diagram, but comes from the conjugacy class of the formal monodromy of the Stokes circle with multiplicity 2, corresponding to the central vertex.
The classes of representations correspond to three different readings of the diagram, as indicated on the following figure:
$\langle 0\rangle_{a}$$\langle 0\rangle_{b}$$\langle 0\rangle_{c}$$\langle 0\rangle_{\infty}$$\bm{\breve{\Theta}}^{VI}_{std}$$\langle\alpha z\rangle_{\infty}$$\langle\beta z\rangle_{\infty}$$\langle\gamma z\rangle_{\infty}$$\langle 0\rangle_{0}$$\bm{\breve{\Theta}}^{VI}_{2}$$\langle\mu z^{2}+\alpha z\rangle_{\infty}$$\langle\mu z^{2}+\beta z\rangle_{\infty}$$\langle\mu z^{2}+\gamma z\rangle_{\infty}$$\langle\lambda z^{2}\rangle_{\infty}$$\bm{\breve{\Theta}}^{VI}_{gen}$
Remark 4.1.
Let us make a few comments about all these Painlevé examples. Notice that, in several cases, the generic tree of the modified irregular class $\bm{\breve{\Theta}}$ breaks part of the symmetry of the equation, for example for Painlevé VI it breaks the symmetry of the 4 regular singular points, and for Painlevé III it breaks the symmetry between the two singularities. This is because the construction depends on the location of the point at infinity, and because to have a minimal number of Stokes circles in $\bm{\breve{\Theta}}$, we made the standard choice of having one of the singularities at infinity. On the other hand, in several cases, the generic tree makes apparent some symmetry that wasn’t directly visible from the fission forest: for example for Painlevé IV the generic tree makes it clear that there is a symmetry between the two Stokes circles at infinity and the two singularities at finite distance, and for degenerate Painlevé III, it makes clear that the forests of two nongeneric classes of representations coincide. The basic reason underlying these observations is that the different combinatorial invariants behave in a nice way with respect to different natural operations on connections: the fission forest for a nonmodified irregular class is invariant under Möbius transformations and twisting by any rank one connection, but not invariant under $\operatorname{SL}_{2}(\mathbb{C})$, while the diagram is invariant under $\operatorname{SL}_{2}(\mathbb{C})$, but not under Möbius transformations changing the location of infinity or twists other than those belonging to the $\operatorname{SL}_{2}(\mathbb{C})$ action. The takeaway from this is that, for any given nonabelian Hodge space, it can be useful to combine all these invariants to understand as many symmetries of the situation as possible.